### v1

# Characterizing the Sensitivity to Individual Bit Flips in Client-Side Operations of the CKKS Scheme

Abstract
Homomorphic Encryption (HE) enables computation on encrypted data without decryption, making it a cornerstone of privacy-preserving computation in untrusted environments. As HE sees growing adoption in sensitive applications–such as secure machine learning and confidential data analysis–ensuring its robustness against errors becomes critical. Faults (e.g., transmission errors, hardware malfunctions, or synchronization failures) can corrupt encrypted data and compromise the integrity of HE operations.
However, the impact of soft errors (such as bit flips) on modern HE schemes remains unexplored. Specifically, the CKKS scheme–one of the most widely used HE schemes for approximate arithmetic–lacks a systematic study of how such errors propagate across its pipeline, particularly under optimizations like the Residue Number System (RNS) and Number Theoretic Transform (NTT).
This work bridges that gap by presenting a theoretical and empirical analysis of CKKS’s fault tolerance under single bit-flip errors. We focus on client-side operations (encoding, encryption, decryption, and decoding) and demonstrate that while the vanilla CKKS scheme exhibits some resilience, performance optimizations (RNS/NTT) introduce significant fragility, amplifying error sensitivity.
By characterizing these failure modes, we lay the groundwork for error-resilient HE designs, ensuring both performance and integrity in privacy-critical applications.
Index Terms:
Homomorphic encryption; resiliency; soft errors; privacy preservation
I Introduction
The need to preserve data privacy becomes critical when processing occurs outside the user’s trusted environment, such as in public clouds or on third-party platforms. In these scenarios, protecting data not only at rest or in transit, but also during processing, is a key challenge.
Homomorphic Encryption (HE) emerges as a promising technique to address this issue. It allows computations to be performed directly on encrypted data without needing to decrypt it first, ensuring confidentiality at all times.
Unlike traditional encryption, which only protects information during storage or transmission, HE enables the secure processing of sensitive data in untrusted environments. For example, it allows a cloud server to calculate statistics on encrypted medical records or train a machine learning model using financial data without revealing its content. In these cases, the server performs the required operations without accessing the plaintext data, preserving its confidentiality throughout the process.
This ability to operate directly on encrypted data is not merely convenient; it is a fundamental requirement for enabling secure computation in infrastructures where full trust cannot be assumed, such as public clouds, third-party platforms, or shared devices. Without such schemes, sensitive applications requiring delegated processing would simply be unfeasible from a privacy standpoint.
Among the most widely used homomorphic schemes is CKKS (Cheon-Kim-Kim-Song), a technique designed for operations on floating-point numbers. This makes it ideal for approximate tasks, such as numerical analysis, signal processing, or machine learning.
To be considered homomorphic, an encryption scheme must be closed under at least addition and multiplication operations on encrypted data. These operations are sufficient to construct any arithmetic circuit, thereby allowing arbitrary computation without needing to decrypt the information. In addition to these fundamental operations, some modern schemes incorporate additional primitives–such as slot rotation or complex conjugation–which, while not increasing the scheme’s expressiveness, are crucial for improving efficiency and facilitating certain computational structures, especially in applications like machine learning.
HE schemes can be divided into three main categories:
  • Partially Homomorphic Encryption (PHE) allows for a single type of operation on encrypted data, typically either additions or multiplications. An example is the traditional RSA[1], which supports the multiplication of encrypted data but not addition.
  • Leveled Homomorphic Encryption (LHE) supports both operations (addition and multiplication), but only for a finite number of steps before the accumulated noise makes decryption returns corrupted data.
  • Fully Homomorphic Encryption (FHE) enables unlimited additions and multiplications, allowing for any arbitrary computation on encrypted data.
Since Craig Gentry’s pioneering proposal [2], numerous studies have optimized the efficiency of HE schemes [3, 4, 5], improving both their software implementations and acceleration through specialized hardware [6, 7, 8].
Due to recent advancements in algorithms and optimizations, it has become feasible to ensure that a service provider cannot interpret, transfer, or access users’ encrypted data, even in the event of a security breach. In cryptographic terms, this is achieved because, with appropriate parameters, breaking a Homomorphic Encryption (HE) scheme becomes a computationally difficult problem. This reinforces confidence in the security of HE and makes its use possible in untrusted environments, such as the public cloud.
Figure 1: FHE used in a typical business application: users send their data encrypted with some homomorphic scheme to the cloud, the server operates using the scheme’s operations and returns a result that it is still encrypted.
As illustrated in Fig. 1, the typical process for an FHE scheme begins with the user encrypting their data locally using their own key before sending it to the server. The server, in turn, can process this data homomorphically; that is, it performs computations directly on the encrypted data. Once processing is complete, the result, which remains encrypted, is returned to the client, who is solely responsible for its decryption.
The CKKS scheme is gaining an increasing interest in machine learning and data analysis applications due to its ability to perform approximate calculations on floating-point numbers. This characteristic is fundamental for efficiently handling large data volumes and training predictive models, allowing sensitive data to be processed securely without needing decryption. Despite its high computational cost, its potential in these fields continues to be an active area of research and development [9, 10, 11].
One of the most significant enhancements to the CKKS scheme is using the Residue Number System (RNS) alongside the Number Theoretic Transform (NTT). This combination has helped establish it as a standard and leads to what is known as the full RNS variant [3]. We will refer to this variant simply as CKKS, while the original version without these optimizations will be termed vanilla CKKS.
It is crucial to recognize that the CKKS scheme is vulnerable to errors caused by single bit-flips, which can compromise both computational accuracy and data integrity. Since bit-flips may occur under various conditions, assessing their impact on CKKS-based applications–particularly in unreliable hardware environments–is essential.
Figure 2 demonstrates the severe impact of a single bit-flip error in a CKKS-encrypted MNIST image (implemented using OpenFHE, a widely adopted FHE library). Despite the ciphertext consisting of millions of bits under standard parameters, flipping just one bit leads to a complete failure in image reconstruction, as evidenced by the corrupted output in Fig. 2(b).
(a) Original
(b) RNS and NTT enabled
Figure 2: Example of the impact of a single bit-flip in CKKS.
These errors can originate in hardware due to phenomena such as interference, memory failures, or even fault injection attacks [12, 13, 14, 15, 16, 17, 18], making bit-flip sensitivity analysis a critical aspect for the practical and reliable implementation of HE.
In this work, we conduct a comprehensive analysis of the CKKS scheme’s sensitivity to single-bit errors in client-side stages, such as encoding, encryption, decryption, and decoding operations. We focus on the impact of RNS (Residue Number System) and NTT (Number Theoretic Transform) optimizations on the scheme’s resilience against these errors, exploring an understudied area in the homomorphic encryption literature.
Unlike other proposals focused on specific applications, our analysis provides a general perspective aimed at establishing a clear framework to enhance the security and robustness of CKKS in diverse contexts. Our main contributions are as follows:
  • We present the first comprehensive analysis of the CKKS scheme’s sensitivity to individual bit-flip errors in a typical client-side pipeline, considering configurations with and without RNS and NTT optimizations.
  • We evaluate the impact of RNS and NTT optimizations on the robustness of the CKKS scheme.
  • We identify specific configurations where the use of certain optimizations introduces a form of intrinsic redundancy, increasing the proportion of bits that remain correct in the presence of errors. This finding is particularly relevant for applications where fault tolerance is a fundamental requirement.
  • We provide a generalized framework that can serve as a foundation for future research aimed at improving the security and resilience of CKKS against hardware failures.
In the rest of the paper, Section II describes in detail the CKKS scheme in its standard version and its full variant with RNS and NTT. Next, Section III presents a theoretical analysis of CKKS’s robustness against bit-flip errors. Then, Section IV details the methodology used in our experimental campaigns. The experimental analysis is developed in Section V, and finally, Section VI presents the conclusions of the work.
II Background and CKKS Overview
In this section, we first introduce the core parameters of the scheme, followed by a conceptual overview alongside a simplified explanation of its key components–making the material accessible even to readers without a specialized background in cryptography.
We then delve into the mathematical foundations required for the analysis presented in Section III. Readers primarily interested in the experimental methodology may proceed directly to Section IV.
II-A General Flow of the CKKS Scheme
CKKS is an approximate homomorphic encryption scheme designed to enable computations on encrypted data while providing explicit control over approximation error. It operates over a polynomial ring of degree $N$ and is based on the Ring Learning With Errors (RLWE) problem, a well-established lattice-based problem that ensures security through the controlled injection of noise into ciphertexts.
CKKS supports the encrypted processing of vectors of real or complex numbers, without ever revealing the original data.
Figure 1 presents the FHE’s process consisting of five main stages:
  1. 1.
[…]
  2. 2.
[…]
    $\boldsymbol{z}=(P(\xi),P(\xi^{3}),\cdots,P(\xi^{2N-1})).$
    A specialized version of the FFT, adapted to the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$, is used to preserve the algebraic structure of the scheme and enable various optimizations.
  3. 3.
    Encryption: Given the plaintext, two encrypted polynomials $(c_{0},c_{1})$ are generated using the public (or secret) key and by adding a small amount of noise. The result $c=(c_{0},c_{1})$ is referred to as the ciphertext.
  4. 4.
    Computation: Homomorphic operations are performed directly on the ciphertexts, including addition, multiplication, rotations, and other transformations compatible with the scheme.
  5. 5.
    Decryption and Decoding: The ciphertext is multiplied with the secret key and summed. The result is then divided by $\Delta$, and a specialized FFT is applied to recover an approximation of the original input vector.
Since the security of CKKS relies on the computational hardness of the RLWE problem, each encryption inherently introduces noise. Although it is small at the beginning, it accumulates during homomorphic operations. In general, each operation increases the noise level in the ciphertext, with homomorphic multiplication contributing the most to this growth.
For this reason, the number of successive multiplications is often used as an indirect estimate of the accumulated error in a ciphertext. If this growth is not properly managed–e.g., through techniques such as bootstrapping–the error can exceed tolerable limits, resulting in incorrect outputs during decryption and decoding.
II-B Parameters controlling CKKS
CKKS operates over a polynomial ring defined as:
$\mathcal{R}_{q}=\mathbb{Z}_{q}[X]/(X^{N}+1),$
where $\mathbb{Z}_{q}$ denotes the integers modulo $q$, that is, the integer values between 0 and $q-1$ (addition and multiplication are performed modulo $q$), $\mathbb{Z}_{q}[X]$ denotes the polynomials with coefficients in $\mathbb{Z}_{q}$, and $/(X^{N}+1)$ indicates that polynomials are considered modulo $X^{N}+1$.
This ring imposes a negacyclic structure that enables, after encoding, the efficient representation of vectors of complex numbers. In CKKS, data is encoded as a polynomial that, when evaluated at specific values, yields each of the encoded data points. In particular, the degree N must be a power of two and serves two critical purposes: it determines the maximum number of complex values that can be encoded simultaneously ($N/2$), and it is essential for maintaining the security of the scheme.
The original (i.e., vanilla) version of the CKKS scheme uses the following core parameters, which are selected to achieve the desired security level, maintain tolerable error growth, and minimize computational cost:
[…]
  • $q_{0}$ is the initial modulus—a single large integer (typically up to 60 bits)–that largely defines the target precision.
  • $\Delta$ is the scaling factor, a constant used to amplify real values before encoding in order to preserve floating-point precision.
  • $L$ is the multiplicative depth, that is, the number of homomorphic multiplications that can be performed before accumulated noise degrades precision.
The total modulus of the scheme, denoted $Q$, and the size of the coefficients in the encoding and encryption polynomials are defined differently depending on the CKKS variant used. In the vanilla version, $Q$ is given by:
$Q=\Delta^{L}\cdot q_{0},$
In CKKS, each multiplication between ciphertexts increases both the error and the magnitude of the encoded values. To control this growth and preserve precision, an operation called rescaling is applied. Rescaling reduces the scale of the encoded values and also decreases the active modulus, which in turn limits the number of consecutive multiplications (with rescaling) that can be performed. This limit is known as the multiplicative depth of the scheme. The parameter $\Delta$ is applied $L$ times–once for each multiplicative level supported by the modulus hierarchy.
II-B1 CKKS full RNS variant
The Residue Number System (RNS) is an alternative representation of integers that expresses a number through its residues modulo several small, pairwise coprime integers. This encoding allows arithmetic operations–such as addition and multiplication–to be performed independently and in parallel on each modulus, reducing computational complexity and significantly improving efficiency. These properties make RNS particularly well-suited for high-performance contexts such as homomorphic encryption.
The use of the Chinese Remainder Theorem (CRT) is essential in RNS, as it ensures that each set of residues uniquely corresponds to a single integer within a defined range. Moreover, CRT provides an efficient way to reconstruct the original integer from its residue representation.
In the RNS-optimized variant of the CKKS scheme, all polynomials involved (i.e., messages, keys, and ciphertexts)are represented as multiple polynomials of the same degree, one for each residue modulo in the RNS base. Additionally, each of these polynomials is transformed using the Number Theoretic Transform (NTT)–a modular analogue of the Fast Fourier Transform (FFT). NTT accelerates polynomial multiplication, which is one of the most computationally intensive operations in the scheme.
These optimizations–RNS and NTT–do not alter the semantics of the encryption scheme but do change its internal implementation. In particular, they have implications for error propagation and sensitivity, topics that will be analyzed in detail in Section III.
In this variant, the total modulus of the scheme is expressed as:
$Q=\prod_{i=0}^{L-1}q_{i},$ (1)
where $q_{0}$ is the initial modulus and $q_{1},\dots,q_{L}$ are additional moduli that in general are selected by the scheme, all of similar bit length. It is worth noting that some libraries provide manual control over the bit-length of these $q_{i}$ moduli, the implementations employed in this work automatically configure these parameters according to the bit length specified for $q_{0}$.
II-C CKKS’ Insights
The CKKS scheme accepts as input a vector of up to $N/2$ floating-point elements, i.e., $\in\mathbb{C}^{N/2}$. This input is mapped into the polynomial ring $\mathbb{Z}[X]/(X^{N}+1)$ through a process known as encoding, producing a plaintext polynomial $m$. Here, $\mathbb{Z}[X]$ denotes the ring of polynomials with integer coefficients, and $\mathbb{Z}[X]/(X^{N}+1)$ is the quotient ring where polynomials are reduced modulo $X^{N}+1$. This implies working with polynomials of degree at most $N-1$ under negative (or negacyclic) convolution. In this context, polynomial reduction is performed via the substitution $X^{N}=-1$, ensuring that results remain within the allowed polynomial space of degree less than $N$.
The resulting plaintext is then encrypted using either the public or secret key, following the RLWE problem, where a small amount of noise is intentionally injected for security. The ring $\mathbb{Z}[X]/(X^{N}+1)$ is preferred in RLWE-based schemes due to its stronger security properties compared to $\mathbb{Z}[X]/(X^{N}-1)$, where the convolution is cyclic and considered more vulnerable to attacks.
Precision
To better understand the parameter selection in CKKS, it is helpful to clarify the concept of precision. In the case of OpenFHE with real-valued inputs only, the scaling parameter $\Delta$ determines how many bits are used to encode the fractional part of the input values, while the difference between the initial modulus $q_{0}$ and $\Delta$ indicates the number of bits allocated to the integer part.
For example, using $q_{0}=60$ bits y $\Delta=50$ bits allows the encrypted representation of real numbers with approximately 10 bits of precision for the integer part and 50 bits for the fractional part.
Keys
Encryption in the CKKS scheme requires either a secret key $sk$ or a public key $pk$. The secret key is generated by sampling a vector $s$ of length $N$ generally from a ternary distribution with Hamming weight $h$, and then setting $sk=(1,s)$.
The public key is constructed by sampling $a_{0}$ from $\mathbb{Z}_{Q}$, i.e., from the set of positive integers less than the modulus $Q$, and an error term $e_{0}$ from a discrete Gaussian distribution with standard deviation $\sigma=3.2$. The public key is then defined as:
$pk=([-a_{0}s+e_{0}]_{Q},a_{0})=(p_{0},p_{1})$
where the notation $[\cdots]_{Q}$ denotes reduction modulo $Q$.
Encoding
The key idea behind encoding in CKKS is to transform the input vector into a polynomial. Different libraries implementing CKKS vary in how they perform this transformation using IFFT/FFT-based methods. In all cases, the standard cyclic transform must be adapted to a negacyclic version to ensure operations remain within the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$ used by the scheme.
The transforms employed by HEAAN 1.0 and OpenFHE 1.2.3 are negacyclic and include scheme-specific optimizations, commonly referred to as specialized FFT/IFFT. These versions incorporate several improvements originally proposed by Bernstein [19, 20]. One of their main advantages is that they allow the transform to be applied directly to vectors of size $\leq N/2$, without expanding them to size $N$, which significantly reduces computational cost.
Given a real-valued input vector of size $n\leq N/2$, the encoding process begins by padding the vector with zeros to reach the next power of two, denoted $n^{\prime}$. A specialized inverse FFT is then applied, producing a complex vector of length $n^{\prime}$. From this, a polynomial of degree $N-1$ is constructed.
First, the gap parameters $=\frac{N/2}{n^{\prime}}$ is computed, and the lower half of the polynomial (coefficients of degree 0 to $N/2-1$) is filled with the real parts of the transformed vector, with assignments spaced at intervals of size gap. Unassigned coefficients are set to zero.
Finally, as is illustrated in Fig. 3, each coefficient is multiplied by $\Delta$, and then rounded to the nearest integer using the coordinate-wise randomized rounding method described in [21].
Figure 3: Encoding scheme in CKKS using HEAAN/OpenFHE. The example shows an input vector with $n=2$ values, represented in a polynomial ring of size $N=8$, where the gap between consecutive coefficients is 2.
Encryption
The encryption of a plaintext $m$ using the secret key $sk=(1,s)$ and modulus $Q$ is given by:
$c_{sk}=([m+a\times s+e]_{Q},[-a]_{Q})=(c_{0},c_{1})$ (2)
where $a$ is a polynomial sampled from a uniform distribution, $e$ is a noise term sampled from a discrete Gaussian distribution, and the notation $[\cdots]_{Q}$ denotes reduction modulo $Q$.
An alternative way is encrypting by using the public key, $pk=(p_{0},p_{1})$. In this case, encryption is defined as:
$c_{pk}=([m+p_{0}\times v+e_{1}]_{Q},[p_{1}\times v+e_{2}]_{Q})=(c_{0},c_{1})$ (3)
where $v$ is an ephemeral (one-time) polynomial sampled from a ternary distribution, and $e_{1}$, $e_{2}$ are error terms sampled similarly to $e$ in the previous procedure.
Decryption
Decryption in both cases (secret key and public key encryption) uses the secret key $sk=(1,s)$ and the encrypted text $(c_{0},c_{1})$ as follows:
$m^{\prime}=[c_{0}+c_{1}\times s]_{Q}$ (4)
Decoding
To decode a plaintext polynomial $m$ with $N$ coefficients, one must reconstruct a complex-valued vector of $n^{\prime}$ elements by reversing the encoding process illustrated in Fig. 3. The $i$-th element ($i\leq N/2$) of the resulting vector is formed by using the real part of the $i$-th coefficient of the plaintext, and the imaginary part from the coefficient at position $i+N/2$.
This complex vector is then divided by $\Delta$, and a specialized FFT is applied to recover the approximate original input.
Due to rounding effects during encoding and decoding, the final result may contain small approximation errors. Corollary 3 in [22] shows that the worst-case distance between the input vector $z$ and the output vector $z^{\prime}$ (after encoding and decoding) at each position is approximately: $N/(\pi\times\Delta)$.
[…]
Security Against Secret Key Attacks
To defend against secret key recovery attacks, OpenFHE performs additional operations before applying the specialized FFT during the decoding process. Specifically, a Gaussian error is injected prior to this step, with its standard deviation determined by the vector being decoded. This mechanism is critical: in our bit-flip experiments–when security analysis is not disabled–the software throws an exception during this stage, resulting in premature termination of execution.
This behavior demonstrates the role of these protective measures in preventing potential side-channel or fault-injection attacks that could compromise the secret key.
II-D RNS and NTT Optimizations
In the full RNS/NTT variant of CKKS [3], the Residue Number System (RNS) is used to split large polynomial coefficients into multiple smaller values, enabling more efficient computations. This approach allows arithmetic in $\mathbb{Z}_{Q}$ to be performed using standard machine word sizes, by working with small moduli that simplify computation.
The magnitude of the modulus $Q$ can range from hundreds to thousands of bits, depending on the use case. However, modern computing systems typically operate with word sizes of 64 bits or less, making direct operations on such large coefficients impractical. To address this, polynomials (keys, plaintexts, ciphertexts) are represented using an equivalent decomposition in the RNS form. Each coefficient is represented modulo a small prime $q_{i}$, to generate $Q$ like Eq. 1
This decomposition enables addition and multiplication in $\mathbb{Z}_{Q}$ using standard machine word operations. All other parameters, as well as the encoding and decoding procedures, remain consistent with the vanilla CKKS variant.
Polynomial multiplications are performed efficiently using the Number Theoretic Transform (NTT), which reduces the computational complexity of the operation from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$. Addition of polynomials and multiplication by a scalar are both $\mathcal{O}(n)$ operations in both the coefficient and evaluation (NTT) representations.
By default, the CKKS scheme assumes that polynomials are represented in the NTT domain (evaluation representation). When certain operations require polynomials in coefficient form, an inverse NTT (iNTT) is applied to transform ciphertexts from the evaluation domain back to the coefficient domain. However, within the scope of this work, as all computations are performed on the client side, this change of representation does not occur.
The procedures for encoding, encryption, decryption, and decoding remain virtually unchanged, except that all operations are now performed using the RNS and NTT representations.
III Error Resilience Analysis
This section presents a theoretical analysis of the error induced by a single bit-flip at different stages of the CKKS scheme. We address both the vanilla and the full RNS variant versions. Our aim is:
  • Deriving analytical bounds in the case without RNS or NTT.
  • Identifying the most sensitive coefficients.
  • Providing an intuitive description of error behavior in practical implementations.
Although a closed-form analytical bound can not be derived for the full RNS variant version due to the complexity of modular representations, a qualitative analysis of the observed patterns is provided.
III-A Error Model for a Single Bit-Flip
We adopt a single bit-flip model, where a single bit is altered. In practice, such faults may be caused by electromagnetic interference, hardware failures, or cosmic rays [12, 15, 23, 13, 24, 17].
Let $p$ be the original value of a coefficient, either in a plaintext polynomial $\boldsymbol{m}$ or in one of the ciphertext polynomials $(c_{0},c_{1})$. We denote this generic polynomial as $\mathbf{p}$. Let’s $p_{j}^{\prime}$ be the result of flipping a bit $j$ of $p$, then the resulting error is:
$e_{j}=p_{j}^{\prime}-p$ (5)
We model this fault as an error vector $\boldsymbol{e}_{i,j}\in\mathbb{R}^{N}$ whose only non-zero component corresponds to bit $j$ of coefficient $i$:
$\mathbf{e}_{i,j}=(0,\dots,0,e_{j},0,\dots,0)$ (6)
This model enables us to track how a single bit-flip propagates through the cryptographic pipeline.
III-B CKKS Encoding via DFT
As discussed in Section II, one of the key components of encoding and decoding in CKKS is FFT. Although CKKS uses a specialized inverse FFT for computational efficiency, this transformation is typically implemented using optimized algorithms that obscure its underlying linear structure.
Instead, we rely on the negacyclic Discrete Fourier Transform (DFT)–an algebraically equivalent transformation explicitly expressed as matrix multiplication. This formulation allows us to:
  1. 1.
    Directly represent how errors propagate through the encoding process.
  2. 2.
    Analyze the isolated effect of a perturbation in a single coefficient.
  3. 3.
    Derive analytical bounds on the magnitude of the error without dealing with the internal optimizations of FFT implementations.
While this choice sacrifices performance, it offers significant pedagogical advantages, allowing for an intuitive understanding of how bit-flip errors are amplified or distributed during encoding.
III-B1 Encoding via Negacyclic DFT
Let $M=2N$ and $\xi=e^{-2\pi i/M}$. To construct a negacyclic Discrete Fourier Transform (DFT), we evaluate it at the odd powers of the roots of the cyclotomic polynomial $\Phi_{M}(X)=X^{N}+1$; that is, at $\xi^{1},\xi^{3},\xi^{5},\dots,\xi^{2N-1}$. We define the Vandermonde matrix where each column corresponds to a different power of these roots of unity:
$W\;=\;\begin{pmatrix}\xi^{1\cdot 0}&\xi^{1\cdot 1}&\cdots&\xi^{1\cdot(N-1)}\\[4.0pt] \xi^{3\cdot 0}&\xi^{3\cdot 1}&\cdots&\xi^{3\cdot(N-1)}\\[2.0pt] \vdots&\vdots&\ddots&\vdots\\[2.0pt] \xi^{(2N-1)\cdot 0}&\xi^{(2N-1)\cdot 1}&\cdots&\xi^{(2N-1)\cdot(N-1)}\end{pmatrix}$ (7)
and we denote its inverse as $W^{-1}$, which is required for the encoding process. Due to the properties of this matrix, its inverse is simply the conjugate transpose (Hermitian transpose) normalized by $N$:
$W^{-1}=\frac{1}{N}W^{\dagger}$
In order to use the DFT and its inverse for encoding to obtain results equivalent to those produced by implementations such as HEAAN and OpenFHE, some preprocessing is required: the input vector must be resized to length $N$ and reordered to satisfy the requirements of those systems. However, these adjustments do not affect the behavior of the error model under analysis and will be omitted for simplicity.
In general, for an input vector of $n\leq N/2$ elements, it is first zero-padded to length $N/2$, and then mirrored with its complex conjugate to reach a total length of $N$. Once the input vector $\boldsymbol{z}\in\mathbb{R}^{N}$ is prepared and reordered, the encoding proceeds as follows:
$\begin{split}\mathrm{Encode}(\boldsymbol{z})&=\mathrm{Round}\bigl{(}W^{-1}(\Delta\boldsymbol{z})\bigr{)}\\ &=\mathrm{Round}\bigl{(}\Delta\,\cdot\,W^{-1}\boldsymbol{z}\bigr{)}=\boldsymbol{m}\in\mathbb{Z}^{N},\end{split}$ (8)
where $\Delta$ is the already defined scaling factor. Due to the Hermitian symmetry of $W$ and $W^{-1}$, and the conjugate extension of the input, the result of $W^{-1}\boldsymbol{z}$ lies entirely in the real subspace.
To recover an approximation of the original input $z$ from $m$, decoding is performed as:
$\mathrm{Decode}(\boldsymbol{m})=\boldsymbol{z}^{\prime}=W\,\bigl{(}\boldsymbol{m}/\Delta\bigr{)}$ (9)
Finally, the output undergoes the same reordering process as in encoding, and the $n$ components of $z^{\prime}$ are extracted. Since the DFT output is complex-valued while OpenFHE supports only real-valued encryption, the real part is extracted to obtain the final approximation of z
III-C Theoretical Estimation of Encoding Error in Vanilla CKKS
This section focuses on the impact of single-bit errors during the encoding phase of the CKKS scheme.
Let us consider the case in which an error $\boldsymbol{e}_{i,j}$ is introduced by flipping a single bit in coefficient $i$ of the plaintext vector $\boldsymbol{m}$, specifically at the $j$-th bit position. The resulting perturbed plaintext is $\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j}$, and decoding proceeds as in Eq. 9:
$\begin{split}\mathbf{z^{\prime\prime}}&=W\times\frac{\mathbf{m^{\prime}}}{\Delta}=W\times\frac{\mathbf{m}+\mathbf{e}_{i,j}}{\Delta}\\ &=W\times\frac{\mathbf{m}}{\Delta}+W\times\frac{\mathbf{e}_{i,j}}{\Delta}\\ &=\mathbf{z^{\prime}}+W\times\frac{\mathbf{e}_{i,j}}{\Delta}\end{split}$
Since a single bit-flip affects only one coefficient, the resulting error is localized. The difference between the decoded vectors $\boldsymbol{z^{\prime}}$ and $\boldsymbol{z^{\prime\prime}}$ is given by:
$\lVert\boldsymbol{z^{\prime}}-\boldsymbol{z^{\prime\prime}}\rVert\propto\lVert W\times\boldsymbol{e}_{i,j}/\Delta\rVert$ (10)
III-C1 Bit-Flip Injection during Encoding
To better understand the analytical expression of the error in Eq. (10), we simulate flipping each individual bit in the plaintext coefficients. Let $j\in\{0,\dots,63\}$ denotes the bit position in a 64-bit unsigned integer. For each iteration, we flip the $j$-th bit of a coefficient, introducing an error with integer magnitude $e_{j}=2^{j}$. This represents the maximum jump in value induced by flipping bit $j$ in a coefficient, which is stored as an unsigned integer–as in the case of plaintext coefficients.
We repeat this process for all coefficients, grouping the results into $N=4$ blocks of 64 iterations each, and use $\Delta=2^{50}$. Let $\boldsymbol{z}^{\prime}$ be the decoded vector without errors, and $\boldsymbol{z}^{\prime\prime}$ the result after applying the bit-flip and decoding. Using Eq. (10), we define:
$L_{2}(i,j)=\bigl{\lVert}Real(W\times\boldsymbol{e}_{i,j}/\Delta)\bigr{\rVert}_{2}$ (11)
where $\mathbf{e}_{i,j}$ is a zero vector of length $N$ with value $e_{j}$ at position $i$.
We also take the real part of the vector before computing the L2 norm, to simulate the encoding of real-valued data, as in OpenFHE, in contrast to HEAAN, which allows complex inputs. This step is important not only to maintain consistency with the backend used in later sections, but also because it exposes a specific behavior that would otherwise not manifest when keeping the complex part.
Figure 4: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding. The horizontal axis shows the bit-flip position $i*64+j$, and the vertical axis shows $L_{2}(i,j)$. Ring dimension $N=4$; scaling factor $\Delta=2^{50}$.
Figure 4 shows the values of $L_{2}(i,j)$ as a function of the bit-flip position. $L2(i,j)$ grows exponentially with $j$ for each coefficient $i$, consistent with the $2^{j}$ dependency. An exception appears at $i=N/2=2$, where the norm remains zero. This occurs because, in the DFT, the coefficient at index $N/2$ contributes only imaginary components, which vanish when taking the real part of the transform.
III-C2 Effect of the Scaling Factor on Bit-Flip Error
To explore how the scaling factor $\Delta$ affects the magnitude of bit-flip errors, we repeat the previous experiment using different values of $\Delta\in\{2^{20},2^{40},2^{50}\}$. For each case, we plot $L_{2}(i,j)$ using a semi-logarithmic scale on the $y$-axis for easier comparison.
Figure 5: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding, for various scaling factors $\Delta$. The horizontal axis indicates the bit position $j$; the vertical axis is in semi-log scale. Ring dimension: $N=4$.
As shown in Fig. 5, increasing $\Delta$ significantly reduces the error magnitude and increases the number of resilient bits. This occurs because the impact of errors becomes noticeable only when the flipped bit corresponds to or exceeds the magnitude of $\Delta$. In other words, the larger the scaling factor, the more initial bits remain unaffected, yielding smaller error norms and higher bit-level fault tolerance.
This empirical observation is analytically confirmed by equation 11, which shows that the error is inversely proportional to the scaling factor. When an error $e_{i,j}$ is divided by $\Delta$, noise components with magnitudes smaller than the bit precision of $\Delta$ are effectively suppressed in the final result. This mathematical relationship explains why larger scaling factors provide enhanced fault resilience: they act as a natural filter that attenuates the impact of small-magnitude bit flips on the decoded output.
However, this resilience comes at the cost of reduced fractional precision in the encoded message (see Section II). The selection of $\Delta$ therefore requires careful consideration of the trade-off between error tolerance and numerical precision, balancing the competing demands of fault resilience and computational accuracy.
III-D Theoretical Estimation of Error during Encryption using Vanilla CKKS
This section analyzes three scenarios involving a single bit-flip in vanilla CKKS:
  1. 1.
    A bit-flip in the plaintext (before encryption).
  2. 2.
    A bit-flip in the ciphertext (in one of the ciphertext polynomials).
  3. 3.
    A bit-flip in the plaintext (before decoding).
Under public-key encryption, a plaintext vector $\mathbf{m}$ is encoded into a polynomial $\boldsymbol{m}(X)$, encrypted as in Eq. 3, and decrypted using Eq. 4.
III-D1 Bit‐Flip in the Plaintext Before Encryption
Suppose a bit-flip introduces an error $\boldsymbol{e}_{i,j}$ in the $i$-th coefficient at bit position $j$, so the plaintext becomes:
$\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j},\quad\boldsymbol{e}_{i,j}=e_{j}\,X^{i-1}$ (12)
Encrypting $\boldsymbol{m}^{\prime}$ using the public key (as in Eq. 3) yields:
$\begin{split}ct^{\prime}&=\bigl{(}[\,\boldsymbol{m}^{\prime}+p_{0}v+e_{1}\,]_{Q},\;[\,p_{1}v+e_{2}\,]_{Q}\bigr{)}\\ &=\bigl{(}[\,\boldsymbol{m}+e_{j}\,X^{i-1}+p_{0}v+e_{1}\,]_{Q},\;c_{1}\bigr{)}=\bigl{(}c_{0}^{\prime},\,c_{1}\bigr{)}.\end{split}$ (13)
The error affects only the ciphertext component $c_{0}^{\prime}$ in coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic. Upon decryption (using Eq. 4):
$\boldsymbol{m}^{\prime\prime}=[\,c_{0}^{\prime}+c_{1}\,s\,]_{Q}=\boldsymbol{m}+e_{j}\,X^{i-1}+(e_{1}+e_{2}\,s)$ (14)
However, the error magnitude grows exponentially with the flipped bit position, as each bit represents a successive power of 2 in the binary representation. Consequently, an exhaustive bit-flip analysis across all coefficient positions, with L2 norm computation for each perturbation, yields the exponential error scaling pattern illustrated in Fig. 4 (SectionIII-C). that the bit-flip impact remains localized to the position $i$ and scales linearly with $e_{j}$.
The error induced by the bit-flip affects only the ciphertext component $c_{0}^{\prime}$ in coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic. Decoding this plaintext will therefore produce an error consistent with that discussed in Section III-C.
III-D2 Bit‐Flip in the Ciphertext
If the bit-flip occurs in the ciphertext, two sub-cases arise:
Flip in $c_{0}$
For a perturbed ciphertext $ct^{\prime}=(c_{0}+e_{j}\,X^{i-1},\;c_{1})$, the behavior mirrors the previous plaintext case: the error remains localized in coefficient $i$ and results in a decoding error proportional to $e_{j}$.
Flip in $c_{1}$
For a perturbed ciphertext $ct^{\prime}=(c_{0},\;c_{1}+e_{j}\,X^{i-1})$, decryption proceeds as:
$\begin{split}\boldsymbol{m}^{\prime\prime}&=[\,c_{0}+(c_{1}+e_{j}\,X^{i-1})\,s\,]_{Q}\\ &=[\,(c_{0}+c_{1}\,s)+(e_{j}\,X^{i-1}\,s)\,]_{Q}\\ &=[\,(\boldsymbol{m}+e_{1}+e_{2}\,s)+(e_{j}\,X^{i-1}\,s)\,]_{Q}\end{split}$
where $e_{j}X^{i-1}\,s$ involves polynomial multiplication. Thus, the error spreads across all $N$ coefficients. Since the secret key $s$ has small coefficients (much less than $q_{0}$), each coefficient in the result incurs an error of magnitude proportional to $e_{j}$, but this error is now distributed. The maximum magnitude is bounded by $|e_{j}|\cdot\|s\|_{\infty}$.
III-D3 Bit‐Flip in the Plaintext Before Decoding
If a bit-flip is introduced in coefficient $i$ of a plaintext $\boldsymbol{m}$, after decryption but before decoding, the effect is analogous to the case analyzed in Section III-C. This follows directly since the plaintext being decoded is equivalent to:
$\boldsymbol{m}^{\prime}=\boldsymbol{m}+e_{j}\,X^{i-1}$ (15)
which is the same structure previously analyzed.
III-E Theoretical Estimation of Error in RNS
CKKS typically uses a Residue Number System (RNS) representation for the modulus $Q=\prod_{k=0}^{L-1}q_{k}$, where $L$ determines the number of available levels (i.e., number of supported homomorphic operations).
Each coefficient $p\in\mathbb{Z}_{Q}$ is represented by its residues:
$r_{k}=p\bmod q_{k}\quad\text{for }k=0,\dots,L$
When applied to all coefficients of a polynomial (from an encoded message or ciphertext), this yields $L$ residue polynomials, commonly referred to as limbs. Reconstruction via the Chinese Remainder Theorem (CRT), as shown in Mohan [25, Chapter 5], is:
$p=\left(\sum_{k=1}^{L}r_{k}\left[\left(\frac{1}{Q_{k}}\right)\bmod q_{k}\right]Q_{k}\right)\text{ mod }Q$ (16)
where
$Q_{k}=Q/q_{k}$
and $\frac{1}{Q_{k}}$ are the multiplicative inverse of $Q_{k}$.
A bit-flip on the $j$ bit affecting the $k$-th residue changes it to $r_{k}^{\prime}=r_{k}+e_{k,j}$, which introduces an error into the reconstructed value:
$p^{\prime}=p+e_{k,j}\,Q_{k}\,\left[\left(\frac{1}{Q_{k}}\right)\bmod q_{k}\right]\pmod{Q}$ (17)
This error can be as large as $Q_{k}$, potentially spanning hundreds or thousands of bits, even though the bit-flip affected only one bit in a single residue.
III-F Theoretical Estimation of Error in NTT
[…]
$\psi^{2}\equiv\xi,\quad\psi^{N}\equiv-1\pmod{q},$
where $\xi$ is a primitive $N$-th root of unity. The NTT matrix is then defined as:
$W_{NTT}=\begin{pmatrix}\psi^{2(0\times 0).0}&\psi^{2(0\times 1).1}&\cdots&\psi^{2(0\times N-1).N-1}\\ \psi^{2(1\times 0).0}&\psi^{2(1\times 1).1}&\cdots&\psi^{2(1\times N-1).N-1}\\ \psi^{2(2\times 0).0}&\psi^{2(2\times 1).1}&\cdots&\psi^{2(2\times N-1).N-1}\\ \vdots&\vdots&\ddots&\vdots\\ \psi^{2(N-1\times 0).0}&\psi^{2(N-1\times 1).1}&\cdots&\psi^{2(N-1\times N-1).N-1}\end{pmatrix}$
We denote its inverse as $W^{-1}_{NTT}$. When a single bit-flip occurs at bit position $j$ within coefficient $i$, the induced perturbation after NTT transformation becomes:
$\mathbf{m}^{\prime}-\mathbf{m}=W_{NTT}\,\mathbf{e}_{i,j}.$
Therefore, the $L_{2}$ norm of the error is:
$\lVert\mathbf{m}^{\prime}-\mathbf{m}\rVert_{2}=\lVert W_{NTT}\,\mathbf{e}_{i,j}\rVert_{2}.$
[…]
Figure 6 shows the values of $L_{1}(i,j)$ for $N=4$ and 64-bit coefficients in the NTT case. The horizontal axis indicates the bit-flip position $i\cdot 64+j$, and the vertical axis shows $L_{2}(i,j)$ on a logarithmic scale.
Figure 6: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding. The horizontal axis shows the modified bit position $i*64+j$; the vertical axis shows the $L_{2}(i,j)$ norm on a logarithmic scale. Ring dimension: $N=4$.
As the Fig. 6 shows, even the flips in the least significant bits cause the norm to exceed four orders of magnitude—producing errors large enough to make correct decoding practically impossible, as will be further illustrated in Section V.
III-G Summary
The following points summarize the theoretical impact of a single bit-flip at different stages of the CKKS encryption pipeline, both in its vanilla and optimized implementations:
  • Bit-flip in the plaintext before encryption: the error remains localized in a single coefficient where the bit-flip occurred, with a magnitude proportional to $e_{j}=2^{j}$, depending on the flipped bit’s position $j$.
  • Bit-flip in the ciphertext ($c_{0}$): identical to the plaintext case—the error remains localized after decryption, with magnitude again proportional to the bit position $j$.
  • Bit-flip in the ciphertext ($c_{1}$): the error is multiplied by the secret key $s$, becoming distributed across all $N$ coefficients; its maximum magnitude is $\mathcal{O}(e_{j}\|s\|_{\infty})$.
  • Bit-flip in the plaintext before decoding: behaves like a localized error in a single coefficient of the encoded polynomial–effectively equivalent to the encoding case, plus minor encryption noise.
  • RNS: a bit-flip in a single limb (residue) can result in a reconstruction error of order $\mathcal{O}(Q_{k})$, where $Q_{k}=Q/q_{k}$ is typically a large value.
[…]
IV Experimental Design
This section describes the experimental design used to quantify the resilience of the CKKS homomorphic encryption scheme under single bit-flip fault injections.
IV-A Test Environment
[…]
  • HEAAN 1.0: a vanilla implementation without RNS or NTT, using arbitrary-precision coefficients managed by the NTL library (a C++ library for number theory and arbitrary-precision arithmetic).
  • OpenFHE 1.2.3: a modular implementation with native support for RNS and NTT, using 64-bit internal coefficients (uint64).
Although more recent versions of HEAAN include RNS and NTT optimizations, version 1.0 allows a cleaner comparison to a vanilla CKKS. Conversely, OpenFHE represents the optimized variant, due to its widespread use in the field of homomorphic encryption [26] and its active support for high-performance execution environments.
Minor modifications were made to both libraries to gain full control over the selection of pseudo-random number generator (PRNG) seeds. This ensured strict reproducibility of experiments and allowed for precise isolation of the effects caused by each individual bit-flip.
In the case of HEAAN 1.0, we have these additional implementation details:
  • The bit-length of each coefficient is not fixed in advance; it depends on internal combinations of $q_{0}$ and $\Delta$, and may vary depending on the PRNG seed used by the scheme.
  • Due to NTL implementation, bits larger than $Q$ are ignored in internal operations, which limits the scope of bit-level analysis.
OpenFHE has the following implementation details:
  • Uses 64-bit integer coefficients (uint64_t) without arbitrary-precision arithmetic.
  • The initial modulus $q_{0}$ can be explicitly configured by the user. This parameter determines the starting bit-level precision of the scheme. The subsequent moduli $q_{i}$ are automatically selected to match the bit length of $q_{0}$, ensuring uniform precision across levels.
  • The total number of moduli depends on the required multiplicative depth of the computation.
  • Public APIs for accessing and modifying coefficients at each pipeline stage (i.e., encoding, encryption, etc.).
  • Enables injection of fixed-length bit-flips (up to 64 bits) at controlled and reproducible positions.
  • Supports 128-bit arithmetic, which–though not used in this work–offers a pathway for future experiments.
IV-B Using OpenFHE as Vanilla CKKS implementation
OpenFHE is significantly more computationally efficient than HEAAN. However, it does not natively offer a mechanism to disable RNS and NTT optimizations. To address this, we devised a procedure to emulate a vanilla CKKS scheme within OpenFHE by explicitly bypassing both RNS and NTT optimizations, despite their built-in support in the library:
  • RNS: We enforced a single-limb representation, which is feasible for depth-zero operations (i.e., no ciphertext multiplications). Additionally, automatic rescaling was disabled and replaced with manual scaling to prevent the introduction of additional limbs.
  • NTT: Prior to injecting a bit-flip, the polynomial is transformed from evaluation space back to coefficient space via inverse NTT (INTT). We flip the desired bit and transform the polynomial back to the evaluation space using NTT. After this hook, the pipeline continue normally.
  • Limitations of disabling NTT: Since the NTT operates modulo $q_{0}$, flipping bits beyond the bit-width of $q_{0}$ causes modular wraparound, which can alter the intended effect of the bit-flip and potentially lead to catastrophic decryption errors. To minimize distortion in the analysis, we set $q_{0}$ to the maximum value supported by the library–60 bits–ensuring that most bit-flips occur within the valid range and reducing unintended side effects from modular arithmetic.
IV-C Internal Error Detection in OpenFHE
OpenFHE includes an internal consistency check that estimates the standard deviation of the imaginary part of the decoded plaintext, designed to detect potential secret key leakage attacks. Flipping high-order bits at different stages of the homomorphic pipeline can result in large deviations, triggering this mechanism and aborting the decoding. To isolate the analysis of numerical resilience, we disable this mechanism.
This methodology enables controlled and reproducible exploration of both vanilla and optimized CKKS configurations, allowing for a rigorous assessment of the impact of single-bit faults at various stages of the homomorphic pipeline.
IV-D Vanilla CKKS over OpenFHE Validation
To ensure that OpenFHE’s vanilla mode faithfully replicates the behavior of HEAAN 1.0, we following these validation steps:
  1. 1.
    Configure OpenFHE in vanilla mode (not using RNS and NTT) and set $q$ and $\Delta$ to match HEAAN’s effective bit precision.
  2. 2.
    Restrict input values in HEAAN to real numbers, in line with OpenFHE’s plaintext constraints.
  3. 3.
    Compare $L_{2}$ norms and error statistics after identical bit-flips in both libraries.
  4. 4.
    Verify consistent sensitivity patterns across low- and high-order bit positions.
IV-E Input Vector Generation and PRNG Control
Since our analysis is not tied to any specific application, the input data consists of real-valued vectors with uniformly distributed random entries. To evaluate the sensitivity of CKKS, we define input ranges using power-of-two boundaries and sample values uniformly within these intervals.
A wider input range affects the signal-to-noise ratio by increasing the difference between the smallest and largest representable values.
Unless otherwise specified, input values are sampled as uniformly distributed real numbers within the interval 0 to 2560256start_ARG 0 end_ARG to start_ARG 256 end_ARG, and represented in double-precision floating-point format. The bounded input range allows us to test multiple $\Delta$, allowing for a more comprehensive analysis of their impact on encoding and rounding errors. By varying $\Delta$ across a wide range without exceeding coefficient bounds, we can better understand the trade-offs between precision and noise growth, as discussed in Section II-B.
To evaluate stability with respect to seed variation, each experiment (i.e., each individual bit-flip) is repeated 2500 times. Each repetition involves 100 different seeds for the scheme PRNG and 25 for the input generator. These seed combinations are fixed and shared across all experiments, ensuring that results could be compared under equivalent input conditions.
Although this number of repetitions may not suffice for exhaustive statistical inference, our tests show very low variability, confirming it is enough for the aims of this study.
IV-F Bit-Flip Injection Procedure
Error injection is performed by flipping a single bit in one encrypted coefficient during each execution. The general workflow for the experiments is as follows:
  1. 1.
    Encode and, if applicable, encrypt the input vector using the selected library.
  2. 2.
    (If required) Apply the inverse NTT (INTT) to bring the plaintext or ciphertext into coefficient space.
  3. 3.
    Flip a single bit within one coefficient (64-bit precision in OpenFHE; arbitrary-precision in HEAAN).
  4. 4.
    (If required) Reapply NTT to return to evaluation space.
  5. 5.
    Proceed with the homomorphic operation pipeline.
  6. 6.
    Decrypt and decode the result.
  7. 7.
    Compute evaluation metrics.
This process is performed once per bit and coefficient across both plaintext and ciphertext scenarios in separate experiments. The entire procedure was repeated for each seed combination.
In OpenFHE, we test configurations both with and without RNS and NTT by dynamically disabling these optimizations at runtime. This allows direct comparisons between the optimized and vanilla implementations, isolating the impact of each optimization technique.
IV-G Evaluation Metrics
Given the original decrypted vector $\mathbf{x}$ (without bit-flips) and the resulting vector $\mathbf{y}$ after error injection and decryption, we quantify the deviation introduced by each bit-flip by two complementary evaluation metrics:
Mean Squared Error (MSE)
The element-wise error is defined as:
$\delta_{i}=y_{i}-x_{i}$
and the MSE is computed as:
$\mathrm{MSE}=\sqrt{\frac{1}{k}\sum_{i=1}^{k}\delta_{i}^{2}}$
Element-wise Relative Error
For each decrypted element, we compute the relative error:
$\epsilon_{i}=\frac{|y_{i}-x_{i}|}{|x_{i}|}$
Results are classified into four categories based on the relative error $\epsilon_{i}$, with each category corresponding to a different tolerance level, which may vary depending on the intended application of FHE:
  1. 1.
    Severe Error ($\epsilon_{i}>10$): the result differs by more than one order of magnitude from the correct value, often rendering it unusable.
  2. 2.
    Significant Error ($0.1<\epsilon_{i}\leq 10$): the result may be unacceptable for precision tasks, but tolerable in coarse-grained applications such as visual representations or robust classification.
  3. 3.
    Moderate Error ($0.01<\epsilon_{i}\leq 0.1$): the deviation is small and potentially acceptable in tasks like statistical analysis or machine learning where high fidelity is not required.
  4. 4.
    Low Error ($\epsilon_{i}\leq 0.01$): the result is nearly indistinguishable from the correct value and is considered acceptably decrypted for most practical purposes.
V Experimental Results
This section presents the results obtained by injecting individual bit-flips during encryption in the CKKS scheme, as described in Section IV, with the goal of evaluating the scheme’s resilience against hardware- or software-induced faults.
Experimental campaigns were carried out using both HEAAN 1.0 and OpenFHE 1.2.3, in their vanilla and optimized configurations (with RNS and NTT).
V-A Vanilla CKKS
V-A1 Evolution of the $L_{2}$ norm under Single Bit‑Flips
To deepen the comparison, we repeated the experiments with HEAAN 1.0, obtaining results that match our theoretical predictions closely. This direct comparison between HEAAN and OpenFHE revealed a marked difference between the vanilla implementation and the fully optimized variant with RNS and NTT: the latter exhibits noticeably greater sensitivity to bit-flips, as evidenced by higher $L_{2}$ error magnitudes across most bit positions.. To isolate the effect of these optimizations, we therefore configured OpenFHE in a simulated vanilla mode by disabling both NTT and RNS, as described in Sec. IV-B.
In order to validate this simulation, we ran both HEAAN and OpenFHE (vanilla) under identical conditions. Because OpenFHE only accepts real inputs, we also restricted HEAAN accordingly to real inputs and, after decryption, considered only the real part of each result. This ensured a homogeneous comparison between the two libraries and confirmed that OpenFHE (vanilla) faithfully reproduces the behavior of HEAAN 1.0, in both trend and magnitude.
Figure 7 shows the case of OpenFHE simulated as CKKS vanilla.
Figure 7: Impact of a single-bit flip during encryption in OpenFHE without RNS or NTT. The horizontal axis indicates the position of the flipped bit within the ciphertext, while the vertical axis shows the resulting $L_{2}$ norm of the decoding error. Parameters: 60‑bit modulus, scaling factor $\Delta=2^{50}$, and ring dimension $N=4$.
We observe that, for the first polynomial $c_{0}$, the behavior mirrors the theoretical analysis in Fig. 4: the least significant bits exhibit higher robustness against bit‑flips, followed by an exponential increase in error magnitude.
In the case of OpenFHE, the expected absolute robustness at coefficient $N/2$ does not hold. During decoding, the coefficient vector prior to the DFT (a specialized FFT) is injected with Gaussian noise affecting both its real and imaginary components. This noise injection is a security measure to thwart key‑recovery attacks.
The introduced Gaussian noise has a standard deviation corresponding to the dispersion of the imaginary part of the vector being transformed, which—by properties of the scheme—also follows a Gaussian distribution. Flipping any bit of any coefficient, including the $N/2$ coefficient, increases this noise standard deviation, especially when higher‑order bits are modified, leading to increased noise variance that propagates across all elements of the vector, significantly altering the final output and negating the robustness at $N/2$.
We experimentally confirmed that, when this noise injection is disabled, the observed behavior aligns with theoretical expectations: the coefficient $N/2$ of $c_{0}$ becomes robust to bit‑flips. Conversely, since HEAAN 1.0 does not perform this noise injection, its sensitivity matches the theory exactly.“‘
V-A2 Impact of the Scaling Factor on Error Categorization
One of the fundamental parameters of the CKKS scheme is the scaling factor $\Delta$. As analyzed in Sec. III, increasing $\Delta$ reduces the relative impact of a single‐bit flip during encoding or encryption and also increases the proportion of bits that remain resilient to such faults.
To investigate this effect in practice, we repeated the previous single‐bit‐flip analysis using four different scaling factors, namely $\Delta\in\{2^{20},2^{30},2^{40},2^{50}\}$. Figure 8 shows the $L_{2}$ norm of the decoding error for a bit flip in the first ciphertext polynomial $c_{0}$ when running OpenFHE in simulated vanilla mode (i.e., with RNS and NTT disabled). For clarity, only results for $c_{0}$ are shown, since the behavior for $c_{1}$ is analogous.
Figure 8: Impact of a single‐bit flip during encryption in OpenFHE without RNS or NTT. The x‑axis indicates the bit position in ciphertext $c_{0}$, while the y‑axis shows the resulting $L_{2}$ norm of the decoding error on a symmetric log scale. Parameters: 60‑bit modulus, scaling factors $\Delta=2^{20},2^{30},2^{40},2^{50}$, and ring dimension $N=8$.
We observe that higher values of $\Delta$ yield lower $L_{2}$ norms after a bit flip, indicating stronger attenuation of induced errors. Furthermore, the point at which the error norm begins its exponential growth shifts in alignment with the magnitude of $\Delta$, confirming that larger scaling factors increase the number of error-resilient bit in each coefficient’s representation.
[…]
For this experiment, we adopted a realistic setting with ring dimension $N=2^{13}=8192$ (i.e., 4096 encrypted slots). For each $\Delta\in\{2^{20},2^{30},2^{40},2^{50}\}$ and each single‐bit flip, we counted how many slots were correctly decrypted and assigned the result to one of the three categories. Figure 9 presents a histogram of the category frequencies for each scaling factor.
[…]
V-A3 Effect of the Effective Representation Range on Error Categorization
We now analyze the impact of the effective dynamic range of the input data representation. We fix $\Delta=2^{20}$, which allows for a wider input range. Recall that the number of bits available for the integer part is approximately given by $q_{0}-\Delta$. The effective dynamic range refers to the ratio between the largest and smallest values that can be represented in fixed-point encoding. In our experiments, this range is controlled indirectly by limiting the number of bits allocated to the fractional part of the input values.
[…]
We observe behavior similar to the impact of the scaling factor, as seen in Fig. 9. As we increase the size of the input space, the percentage of correctly decrypted elements increases while the percentage of incorrectly decrypted elements decreases. If the size of input elements grows while maintaining a fixed scaling factor, this allows representing values with greater relative precision compared to the existing noise.
V-A4 Effect of reducing the number of encrypted elements
As noted previously, CKKS can encode up to $N/2$ slots, where $N$ is the ring degree (a power of two). When we encode fewer slots—still a power of two—we observe that additional coefficients remain unaffected by single‐bit flips. Figure 11 shows results for $N=16$ with only $N/4=4$ encrypted slots, allowing us to inspect a larger set of coefficients and evaluate their sensitivity.
Figure 11: Impact of a single‐bit flip during encryption in OpenFHE without RNS or NTT. The horizontal axis indicates the bit position in the ciphertext, while the vertical axis shows the resulting $L_{2}$ norm of the decoding error. Parameters: 60‑bit modulus, scaling factor $\Delta=2^{20}$, ring dimension $N=16$, and 4 encrypted slots.
The results reveal that odd‐indexed coefficients of polynomial $c_{0}$ remain unchanged under bit‐flips, whereas even‐indexed coefficients exhibit the same sensitivity patterns observed earlier.
This phenomenon stems from CKKS’s decoding process in both OpenFHE and HEAAN. As explained in Sec. II, the ratio between the maximum number of slots ($N/2$) and the number actually used defines an interval—termed the gap—between the coefficients participating in the inverse transform. During the “special” FFT, only coefficients whose indices are multiples of this gap are considered and scaled in the input vector (see Sec. I for details).
Consequently, when encoding $N/2$ slots, all $N$ coefficients of $c_{0}$ are sensitive to bit‐flips; encoding $N/4$ slots reduces the number of sensitive coefficients of $c_{0}$ to $N/2$; and so on. In the case of $c_{1}$, no matter the number of slots used, all coefficients are sensitive to bit changes. Experimental campaigns confirm that our OpenFHE vanilla simulation (with RNS and NTT disabled) matches the expected behavior.
V-B Resilience under an intermediate RNS mode without NTT
We next conducted experimental campaigns to assess CKKS’s sensitivity when using only one of the numeric optimizations—either NTT or RNS. I n both pure-NTT and pure-RNS configurations, any single‐bit flip caused decryption errors of unacceptable magnitude. As shown in our theoretical analysis, bit flips under RNS produce errors several orders of magnitude larger than under NTT, while NTT alone still yields extremely large errors that preclude practical precision.
An interesting case emerges when CKKS is run with RNS only and the number of encrypted slots is halved or more. Figure 12 reports results for ring dimension $N=16$ with only 444 encrypted slots (i.e., one power‐of‐two below the maximum), scaling factor $\Delta=2^{20}$, and 8‐bit input values.
[…]
Polynomial $c_{1}$ exhibits no robust bits at all: every bit flip causes uniform decoding failure, so its error‐categorization histogram is flat and standard deviation is zero.
These results demonstrate that running CKKS with RNS only—together with a non‐maximal but non‐zero gap—can concentrate robustness on a subset of coefficients. Such a configuration suggests a trade‐off between ring utilization and fault resilience. Future work may explore tailored RNS and NTT parameterizations to minimize sensitivity to bit flips while preserving sufficient slot capacity, laying the groundwork for error‐reduction strategies in homomorphic encryption schemes that balance precision and fault tolerance.“‘
V-C Summary and Comparison
We provides a high‑level comparison of the different CKKS configurations evaluated in this work. For each scenario, we highlight the key resilience pattern and any notable observations.
  • Vanilla CKKS:
    • Least‐significant bits exhibit robustness; error grows exponentially for higher bits.
    • Coefficient $N/2$ robust in HEAAN 1.0 but not in OpenFHE vanilla (due to Gaussian noise injection).
  • Varying $\Delta$:
    • Larger $\Delta$ delays the onset of exponential error growth to more significant bits.
    • $\Delta\geq 2^{30}$ yields $\mathbf{Tail}$ outcomes in over 80% of trials.
  • Effective Representation Range:
    • Increasing integer‐part range raises $\mathbf{Tail}$ frequency and reduces $\mathbf{Head}$ frequency.
    • Behavior parallels that of increasing $\Delta$.
  • Reduced Slots:
    • Encoding fewer slots reduce the number of sensitive coefficients in $c_{0}$.
  • RNS without NTT:
    • Some coefficients of $c_{0}$ remain robust; all flips in $c_{1}$ fail.
    • Demonstrates a “gap”‐driven concentration of robustness when using RNS only.
This concise overview highlights how each parameter adjustment—from vanilla through optimized variants—modulates CKKS’s fault tolerance, guiding practical parameter selection under two‐column constraints.
VI Conclusions
This paper has presented a comprehensive evaluation of CKKS’s sensitivity to single‐bit flips, emphasizing the influence of RNS and NTT optimizations on error resilience. While the vanilla CKKS implementation exhibits substantial innate tolerance to individual bit‐flip faults, the optimized variants—with RNS or NTT enabled—display markedly increased vulnerability, potentially undermining the correctness of decrypted outputs.
Our findings indicate that, despite its performance drawbacks, vanilla CKKS offers a robustness advantage by minimizing the impact of isolated bit errors compared to its optimized counterparts. Furthermore, when employing large ring dimensions for security, encoding fewer slots (powers of two below the maximum) introduces redundancy that renders a subset of coefficients entirely insensitive to bit flips. This “gap”‐driven effect persists even in optimized configurations, highlighting a practical mechanism for enhancing resilience without altering core codec operations.
Overall, these results deepen our understanding of fault tolerance in homomorphic encryption schemes and provide actionable guidance for balancing throughput, security, and error resilience when selecting CKKS parameters.
Future Work
Promising directions include studying error tolerance in application‐specific domains—such as deep neural networks for image processing, where limited accuracy loss is acceptable—and conducting a detailed investigation of error propagation under multiple concurrent bit flips. Such analyses will further illuminate strategies for error mitigation in practical homomorphic computing scenarios.
Acknowledgments
First, I would like to thank Pradip Bose, whose inspiration to explore fault tolerance in homomorphic encryption significantly shaped this work. I also appreciate the open‐source homomorphic encryption community and the developers of the OpenFHE and SEAL libraries, whose tools and documentation greatly facilitated our experiments. In particular, I extend my gratitude to Ahmad Al Badawi for his help setting up OpenFHE correctly and for his clear explanations and references on CKKS encoding and decoding.

### v3

@@
 Faults (e.g., transmission errors, hardware malfunctions, or synchronization failures) can corrupt encrypted data and compromise the integrity of HE operations.
-However, the impact of soft errors (such as bit flips) on modern HE schemes remains unexplored.
+However, the impact of soft errors (such as bit-flips) on modern HE schemes remains unexplored.
 Specifically, the CKKS scheme–one of the most widely used HE schemes for approximate arithmetic–lacks a systematic study of how such errors propagate across its pipeline, particularly under optimizations like the Residue Number System (RNS) and Number Theoretic Transform (NTT).
-This work bridges that gap by presenting a theoretical and empirical analysis of CKKS’s fault tolerance under single bit-flip errors.
-We focus on client-side operations (encoding, encryption, decryption, and decoding) and demonstrate that while the vanilla CKKS scheme exhibits some resilience, performance optimizations (RNS/NTT) introduce significant fragility, amplifying error sensitivity.
-By characterizing these failure modes, we lay the groundwork for error-resilient HE designs, ensuring both performance and integrity in privacy-critical applications.
+This work bridges this gap by presenting a theoretical and empirical analysis of CKKS’s fault tolerance behavior under bit-flip errors.
+We focus on client-side operations (encoding, encryption, decryption, and decoding) and demonstrate that while the vanilla CKKS scheme exhibits some resilience, performance optimizations like RNS and NTT exacerbate the error sensitivity of the scheme.
+By characterizing these failure modes, we lay the groundwork for developing HE hardware and software systems that are resilient to errors, ensuring both performance and integrity in privacy-critical applications.
 Index Terms:
@@
 I Introduction
-The need to preserve data privacy becomes critical when processing occurs outside the user’s trusted environment, such as in public clouds or on third-party platforms.
+The need to preserve data privacy becomes critical when processing occurs outside the client’s trusted environment, such as in public clouds or on third-party platforms.
 In these scenarios, protecting data not only at rest or in transit, but also during processing, is a key challenge.
-Homomorphic Encryption (HE) emerges as a promising technique to address this issue.
-It allows computations to be performed directly on encrypted data without needing to decrypt it first, ensuring confidentiality at all times.
-Unlike traditional encryption, which only protects information during storage or transmission, HE enables the secure processing of sensitive data in untrusted environments.
+Homomorphic Encryption (HE) emerges as a promising technique to address this issue by enabling computation on encrypted data (i.e., without requiring decryption), thereby ensuring confidentiality at all times.
 For example, it allows a cloud server to calculate statistics on encrypted medical records or train a machine learning model using financial data without revealing its content.
@@
 This ability to operate directly on encrypted data is not merely convenient; it is a fundamental requirement for enabling secure computation in infrastructures where full trust cannot be assumed, such as public clouds, third-party platforms, or shared devices.
-Without such schemes, sensitive applications requiring delegated processing would simply be unfeasible from a privacy standpoint.
-Among the most widely used homomorphic schemes is CKKS (Cheon-Kim-Kim-Song), a technique designed for operations on floating-point numbers.
-This makes it ideal for approximate tasks, such as numerical analysis, signal processing, or machine learning.
-To be considered homomorphic, an encryption scheme must be closed under at least addition and multiplication operations on encrypted data.
-These operations are sufficient to construct any arithmetic circuit, thereby allowing arbitrary computation without needing to decrypt the information.
-In addition to these fundamental operations, some modern schemes incorporate additional primitives–such as slot rotation or complex conjugation–which, while not increasing the scheme’s expressiveness, are crucial for improving efficiency and facilitating certain computational structures, especially in applications like machine learning.
+Without such schemes, sensitive applications requiring delegated processing would be unfeasible from a privacy standpoint.
+Among the most widely used homomorphic schemes is CKKS (Cheon-Kim-Kim-Song) [1] due to its fixed-point arithmetic support.
+This makes it ideal to handle real number computation in tasks such as numerical analysis, signal processing, or machine learning.
+Figure 1: Typical HE use case: a client sends encrypted data to the cloud, the cloud operates on the encrypted data and returns the results.
+Data and results remain encrypted all the time.
+[^1] Adapted from https://openmined.org/blog/from-fully-homomorphic-encryption-to-silicon/
+An encryption scheme should support encrypted computation of at least addition and multiplication operations in order to be considered homomorphic.
+These operations are sufficient to construct any arithmetic circuit, thereby allowing arbitrary computation without decrypting the information.
+In addition to these fundamental operations, some modern schemes incorporate additional primitives–such as elements rotation or complex conjugation–which, while not increasing the scheme’s expressiveness, are crucial for improving efficiency and facilitating certain computational structures, especially in applications like machine learning.
 HE schemes can be divided into three main categories:
-  • Partially Homomorphic Encryption (PHE) allows for a single type of operation on encrypted data, typically either additions or multiplications.
-An example is the traditional RSA[1], which supports the multiplication of encrypted data but not addition.
-  • Leveled Homomorphic Encryption (LHE) supports both operations (addition and multiplication), but only for a finite number of steps before the accumulated noise makes decryption returns corrupted data.
-  • Fully Homomorphic Encryption (FHE) enables unlimited additions and multiplications, allowing for any arbitrary computation on encrypted data.
-Since Craig Gentry’s pioneering proposal [2], numerous studies have optimized the efficiency of HE schemes [3, 4, 5], improving both their software implementations and acceleration through specialized hardware [6, 7, 8].
-Due to recent advancements in algorithms and optimizations, it has become feasible to ensure that a service provider cannot interpret, transfer, or access users’ encrypted data, even in the event of a security breach.
-In cryptographic terms, this is achieved because, with appropriate parameters, breaking a Homomorphic Encryption (HE) scheme becomes a computationally difficult problem.
-This reinforces confidence in the security of HE and makes its use possible in untrusted environments, such as the public cloud.
-Figure 1: FHE used in a typical business application: users send their data encrypted with some homomorphic scheme to the cloud, the server operates using the scheme’s operations and returns a result that it is still encrypted.
-As illustrated in Fig.
-1, the typical process for an FHE scheme begins with the user encrypting their data locally using their own key before sending it to the server.
-The server, in turn, can process this data homomorphically; that is, it performs computations directly on the encrypted data.
+  • Partial Homomorphic Encryption (PHE) supports a single type of operation on encrypted data, either addition or multiplication.
+An example is The RSA (Rivest–Shamir–Adleman) cryptosystem[2], which supports multiplication of encrypted data but not addition.
+  • Leveled Homomorphic Encryption (LHE) supports both operations (addition and multiplication), but only for a limited number of steps before the accumulated noise makes correct decryption infeasible.
+  • Fully Homomorphic Encryption (FHE) enables unlimited additions and multiplications, enabling arbitrary computation on encrypted data.
+As illustrated in Figure 1, the typical process for an HE scheme begins with the client encrypting data locally using a secret key before sending it to the cloud server.
+The server process this data homomorphically; that is, it performs computations directly on the encrypted data.
 Once processing is complete, the result, which remains encrypted, is returned to the client, who is solely responsible for its decryption.
-The CKKS scheme is gaining an increasing interest in machine learning and data analysis applications due to its ability to perform approximate calculations on floating-point numbers.
+The CKKS scheme is gaining increasing interest in machine learning and data analysis applications due to its ability to perform approximate calculations on floating-point numbers.
 This characteristic is fundamental for efficiently handling large data volumes and training predictive models, allowing sensitive data to be processed securely without needing decryption.
-Despite its high computational cost, its potential in these fields continues to be an active area of research and development [9, 10, 11].
+Despite its high computational cost, its potential in these fields is still an active area of research and development [3, 4, 5].
 One of the most significant enhancements to the CKKS scheme is using the Residue Number System (RNS) alongside the Number Theoretic Transform (NTT).
 This combination has helped establish it as a standard and leads to what is known as the full RNS variant [6].
-We will refer to this variant simply as CKKS, while the original version without these optimizations will be termed vanilla CKKS.
+Throughout this paper, we will refer to this variant simply as CKKS, while the original version without these optimizations will be termed vanilla CKKS.
 It is crucial to recognize that the CKKS scheme is vulnerable to errors caused by single bit-flips, which can compromise both computational accuracy and data integrity.
-Since bit-flips may occur under various conditions, assessing their impact on CKKS-based applications–particularly in unreliable hardware environments–is essential.
-Figure 2 demonstrates the severe impact of a single bit-flip error in a CKKS-encrypted MNIST image (implemented using OpenFHE, a widely adopted FHE library).
-Despite the ciphertext consisting of millions of bits under standard parameters, flipping just one bit leads to a complete failure in image reconstruction, as evidenced by the corrupted output in Fig.
-2(b).
+Since bit-flips may occur under various conditions, assessing their impact on CKKS-based applications–particularly in unreliable hardware environments–is crucial.
+Figure 2 illustrates a scenario in which an image from the MNIST dataset (Figure 2(a)) is processed through the CKKS pipeline.
+Figure 2(b) shows an entirely distorted image recovered when a single bit error is introduced during the encoding phase using RNS and NTT.
 (a) Original
@@
 These errors can originate in hardware due to phenomena such as interference, memory failures, or even fault injection attacks [7, 8, 9, 10, 11, 12, 13], making bit-flip sensitivity analysis a critical aspect for the practical and reliable implementation of HE.
-In this work, we conduct a comprehensive analysis of the CKKS scheme’s sensitivity to single-bit errors in client-side stages, such as encoding, encryption, decryption, and decoding operations.
-We focus on the impact of RNS (Residue Number System) and NTT (Number Theoretic Transform) optimizations on the scheme’s resilience against these errors, exploring an understudied area in the homomorphic encryption literature.
-Unlike other proposals focused on specific applications, our analysis provides a general perspective aimed at establishing a clear framework to enhance the security and robustness of CKKS in diverse contexts.
+In this work, we conduct a comprehensive analysis of the CKKS scheme’s sensitivity to single bit errors in client-side stages, such as encoding, encryption, decryption, and decoding operations.
+We focus on the impact of RNS (Residue Number System) and NTT (Number Theoretic Transform) optimizations on the scheme’s resilience against these errors.
+To the best of our knowledge, this paper is the first study that explores this aspect of HE.
+Our analysis provides a general framework aimed at enhancing the security and robustness of CKKS in diverse contexts.
 Our main contributions are as follows:
@@
 This finding is particularly relevant for applications where fault tolerance is a fundamental requirement.
-  • We provide a generalized framework that can serve as a foundation for future research aimed at improving the security and resilience of CKKS against hardware failures.
-In the rest of the paper, Section II describes in detail the CKKS scheme in its standard version and its full variant with RNS and NTT.
-Next, Section III presents a theoretical analysis of CKKS’s robustness against bit-flip errors.
-Then, Section IV details the methodology used in our experimental campaigns.
-The experimental analysis is developed in Section V, and finally, Section VI presents the conclusions of the work.
+  • We provide a framework that can serve as a foundation for future research aimed at improving the security and resilience of CKKS against hardware failures.
+Section II describes the CKKS scheme in its vanilla and full variants.
+Section III presents a theoretical analysis of CKKS’s robustness against bit-flip errors.
+Section IV details the methodology used in our experimental campaigns.
+Section V develops the experimental analysis and discusses our findings.
+Finally, Section VI presents our conclusions.
 II Background and CKKS Overview
-In this section, we first introduce the core parameters of the scheme, followed by a conceptual overview alongside a simplified explanation of its key components–making the material accessible even to readers without a specialized background in cryptography.
-We then delve into the mathematical foundations required for the analysis presented in Section III. Readers primarily interested in the experimental methodology may proceed directly to Section IV.
+In this section, we first introduce the core parameters of the CKKS scheme, followed by a conceptual overview alongside a simplified explanation of its key components–making the material accessible even to readers without a specialized background in cryptography.
+We then delve into the mathematical foundations required for the analysis presented in Section III.
 II-A General Flow of the CKKS Scheme
 CKKS is an approximate homomorphic encryption scheme designed to enable computations on encrypted data while providing explicit control over approximation error.
-It operates over a polynomial ring of degree $N$ and is based on the Ring Learning With Errors (RLWE) problem, a well-established lattice-based problem that ensures security through the controlled injection of noise into ciphertexts.
-CKKS supports the encrypted processing of vectors of real or complex numbers, without ever revealing the original data.
-Figure 1 presents the FHE’s process consisting of five main stages:
+It operates over a polynomial ring of degree $N$ based on the Ring Learning With Errors (RLWE) problem, a lattice-based problem that ensures security through the controlled injection of noise into ciphertexts.
+Figure 1 shows the processing stages:
   1. 1.
-[…]
+    Key Generation: The secret key and public key are generated as degree-$N$ polynomials, created through random sampling and error injection.
   2. 2.
-[…]
-    $\boldsymbol{z}=(P(\xi),P(\xi^{3}),\cdots,P(\xi^{2N-1})).$
+    Encoding: The input vector is transformed using a specialized inverse Fast Fourier Transform (IFFT) into a degree-$N$ polynomial, constructed such that its evaluations at certain roots of unity approximately reproduce the original values.
+The resulting polynomial’s coefficients are then scaled by $\Delta$ and rounded to integers, producing the plaintext.
+Formally, given a polynomial $P$ and a vector $z\in\mathbb{R}^{N}$, the following relation holds:
+    $\boldsymbol{z}\approx(P(\xi),P(\xi^{3}),\cdots,P(\xi^{2N-1})).$
     A specialized version of the FFT, adapted to the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$, is used to preserve the algebraic structure of the scheme and enable various optimizations.
@@
   5. 5.
-    Decryption and Decoding: The ciphertext is multiplied with the secret key and summed.
-The result is then divided by $\Delta$, and a specialized FFT is applied to recover an approximation of the original input vector.
+    Decryption and Decoding: One component of the ciphertext is multiplied by the secret key and added to the other component.
+The result is a plaintext, which is divided by $\Delta$, and then a specialized FFT is applied to recover an approximation of the original input vector.
 Since the security of CKKS relies on the computational hardness of the RLWE problem, each encryption inherently introduces noise.
 Although it is small at the beginning, it accumulates during homomorphic operations.
-In general, each operation increases the noise level in the ciphertext, with homomorphic multiplication contributing the most to this growth.
+Each operation increases the noise level in the ciphertext, with homomorphic multiplication contributing the most to this growth.
 For this reason, the number of successive multiplications is often used as an indirect estimate of the accumulated error in a ciphertext.
-If this growth is not properly managed–e.g., through techniques such as bootstrapping–the error can exceed tolerable limits, resulting in incorrect outputs during decryption and decoding.
+If this growth is not properly managed–e.g., through techniques such as bootstrapping [14]–the error can exceed tolerable limits, resulting in incorrect outputs during decryption and decoding.
 II-B Parameters controlling CKKS
@@
 $\mathcal{R}_{q}=\mathbb{Z}_{q}[X]/(X^{N}+1),$
-where $\mathbb{Z}_{q}$ denotes the integers modulo $q$, that is, the integer values between 0 and $q-1$ (addition and multiplication are performed modulo $q$), $\mathbb{Z}_{q}[X]$ denotes the polynomials with coefficients in $\mathbb{Z}_{q}$, and $/(X^{N}+1)$ indicates that polynomials are considered modulo $X^{N}+1$.
+where $\mathbb{Z}_{q}$ denotes the integers modulo $q$ (i.e., the integer values between 0 and q1q-1, with addition and multiplication performed modulo $q$)); $\mathbb{Z}_{q}[X]$ denotes the polynomials with coefficients in $\mathbb{Z}_{q}$; and $(X^{N}+1)$ indicates that polynomials are considered modulo $X^{N}+1$.
 This ring imposes a negacyclic structure that enables, after encoding, the efficient representation of vectors of complex numbers.
 In CKKS, data is encoded as a polynomial that, when evaluated at specific values, yields each of the encoded data points.
-In particular, the degree N must be a power of two and serves two critical purposes: it determines the maximum number of complex values that can be encoded simultaneously ($N/2$), and it is essential for maintaining the security of the scheme.
-The original (i.e., vanilla) version of the CKKS scheme uses the following core parameters, which are selected to achieve the desired security level, maintain tolerable error growth, and minimize computational cost:
-[…]
-  • $q_{0}$ is the initial modulus—a single large integer (typically up to 60 bits)–that largely defines the target precision.
-  • $\Delta$ is the scaling factor, a constant used to amplify real values before encoding in order to preserve floating-point precision.
-  • $L$ is the multiplicative depth, that is, the number of homomorphic multiplications that can be performed before accumulated noise degrades precision.
+In particular, the degree $N$ must be a power of two and serves two critical purposes: it determines the maximum number of complex values that can be encoded simultaneously, and enforces the security of the scheme.
+The original (vanilla) version of the CKKS scheme uses the following core parameters, which are selected to achieve the desired security level, maintain tolerable error growth, and minimize computational cost:
+  • NN is the ring degree, typically in the range $2^{12}-2^{16}$, and it determines both the dimension of the encoding space and the computational cost.
+  • $q_{0}$ is the initial modulus–a single large integer (typically up to 60 bits)–that determines the target precision.
+  • $\Delta$ is the scaling factor, a constant used to amplify real values before encoding to preserve floating-point precision.
+  • $L$ is the multiplicative depth; i.e., the number of homomorphic multiplications that can be performed before accumulated noise degrades precision.
 The total modulus of the scheme, denoted $Q$, and the size of the coefficients in the encoding and encryption polynomials are defined differently depending on the CKKS variant used.
-In the vanilla version, $Q$ is given by:
-$Q=\Delta^{L}\cdot q_{0},$
+In the vanilla version, $Q=\Delta^{L}\cdot q_{0}$.
 In CKKS, each multiplication between ciphertexts increases both the error and the magnitude of the encoded values.
@@
 This limit is known as the multiplicative depth of the scheme.
-The parameter $\Delta$ is applied $L$ times–once for each multiplicative level supported by the modulus hierarchy.
-II-B1 CKKS full RNS variant
-The Residue Number System (RNS) is an alternative representation of integers that expresses a number through its residues modulo several small, pairwise coprime integers.
-This encoding allows arithmetic operations–such as addition and multiplication–to be performed independently and in parallel on each modulus, reducing computational complexity and significantly improving efficiency.
+The factor $\Delta$ appears once at each of the $L$ multiplicative levels of the modulus hierarchy, which explains the term $\Delta^{L}$ in the expression for $Q$.
+II-C Full CKKS Variant
+The magnitude of the modulus $Q$ (and consequently, the size of the polynomial coefficients) can range from hundreds to thousands of bits, depending on the use case.
+However, modern computing systems typically operate with word sizes of 64 bits or less, making direct operations on such large coefficients impractical.
+The Residue Number System (RNS) enables the representation of integer values using their remainders with respect to a set of pairwise coprime moduli.
+This allows arithmetic operations–such as addition and multiplication–to be performed using 64-bit word size, reducing computational complexity and significantly improving efficiency.
 These properties make RNS particularly well-suited for high-performance contexts such as homomorphic encryption.
 The use of the Chinese Remainder Theorem (CRT) is essential in RNS, as it ensures that each set of residues uniquely corresponds to a single integer within a defined range.
-Moreover, CRT provides an efficient way to reconstruct the original integer from its residue representation.
-In the RNS-optimized variant of the CKKS scheme, all polynomials involved (i.e., messages, keys, and ciphertexts)are represented as multiple polynomials of the same degree, one for each residue modulo in the RNS base.
-Additionally, each of these polynomials is transformed using the Number Theoretic Transform (NTT)–a modular analogue of the Fast Fourier Transform (FFT).
-NTT accelerates polynomial multiplication, which is one of the most computationally intensive operations in the scheme.
-These optimizations–RNS and NTT–do not alter the semantics of the encryption scheme but do change its internal implementation.
+Moreover, CRT provides an efficient procedure to reconstruct the original integer from its residue representation.
+Once mapped to the Residue Number System (RNS), each polynomial is further transformed via the Number Theoretic Transform (NTT)–the modular analogue of the Fast Fourier Transform (FFT).
+This reduces the complexity of polynomial multiplication (one of the most computationally demanding operations in the scheme) from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$.
+By default, the CKKS scheme assumes that polynomials are represented in the evaluation domain (NTT).
+When certain operations require polynomials to be in the coefficient domain, an inverse NTT (iNTT) is applied to transform the polynomial from the evaluation domain back to the coefficient domain.
+These optimizations (RNS and NTT) do not alter the semantics of the encryption scheme but do change its internal implementation.
 In particular, they have implications for error propagation and sensitivity, topics that will be analyzed in detail in Section III.
-In this variant, the total modulus of the scheme is expressed as:
+In the full CKKS variant, the total modulus of the scheme is expressed as:
 $Q=\prod_{i=0}^{L-1}q_{i},$ (1)
-where $q_{0}$ is the initial modulus and $q_{1},\dots,q_{L}$ are additional moduli that in general are selected by the scheme, all of similar bit length.
-It is worth noting that some libraries provide manual control over the bit-length of these $q_{i}$ moduli, the implementations employed in this work automatically configure these parameters according to the bit length specified for $q_{0}$.
-II-C CKKS’ Insights
-The CKKS scheme accepts as input a vector of up to $N/2$ floating-point elements, i.e., $\in\mathbb{C}^{N/2}$.
-This input is mapped into the polynomial ring $\mathbb{Z}[X]/(X^{N}+1)$ through a process known as encoding, producing a plaintext polynomial $m$.
-Here, $\mathbb{Z}[X]$ denotes the ring of polynomials with integer coefficients, and $\mathbb{Z}[X]/(X^{N}+1)$ is the quotient ring where polynomials are reduced modulo $X^{N}+1$.
-This implies working with polynomials of degree at most $N-1$ under negative (or negacyclic) convolution.
-In this context, polynomial reduction is performed via the substitution $X^{N}=-1$, ensuring that results remain within the allowed polynomial space of degree less than $N$.
+where $q_{0}$ is the initial modulus and $q_{1},\dots,q_{L-1}$ are additional moduli that in general are selected by the scheme, all of similar bit length.
+It is worth noting that some libraries provide manual control over the bit-length of these $q_{i}$ moduli; the implementations employed in this work automatically configure these parameters according to the bit length specified for $q_{0}$.
+II-D Encoding/Encryption/Decryption/Decoding
+The CKKS scheme operates on complex vectors of size up to $N/2$.
+This input is encoded into the polynomial ring $\mathbb{Z}[X]/(X^{N}+1)$, producing a plaintext polynomial $m$.
 The resulting plaintext is then encrypted using either the public or secret key, following the RLWE problem, where a small amount of noise is intentionally injected for security.
@@
 To better understand the parameter selection in CKKS, it is helpful to clarify the concept of precision.
-In the case of OpenFHE with real-valued inputs only, the scaling parameter $\Delta$ determines how many bits are used to encode the fractional part of the input values, while the difference between the initial modulus $q_{0}$ and $\Delta$ indicates the number of bits allocated to the integer part.
-For example, using $q_{0}=60$ bits y $\Delta=50$ bits allows the encrypted representation of real numbers with approximately 10 bits of precision for the integer part and 50 bits for the fractional part.
+Some CKKS implementations use real-valued inputs only, the scaling parameter $\Delta$ determines how many bits are used to encode the fractional part of the input values, and the difference between the initial modulus $q_{0}$ and $\Delta$ indicates the number of bits allocated to the integer part.
+For example, using $q_{0}=60$ bits and $\Delta=50$ bits allows the encrypted representation of real numbers with approximately 10 bits of precision for the integer part and 50 bits for the fractional part.
 Keys
 Encryption in the CKKS scheme requires either a secret key $sk$ or a public key $pk$.
-The secret key is generated by sampling a vector $s$ of length $N$ generally from a ternary distribution with Hamming weight $h$, and then setting $sk=(1,s)$.
-The public key is constructed by sampling $a_{0}$ from $\mathbb{Z}_{Q}$, i.e., from the set of positive integers less than the modulus $Q$, and an error term $e_{0}$ from a discrete Gaussian distribution with standard deviation $\sigma=3.2$.
-The public key is then defined as:
+The secret key is generated as $sk=(1,s)$ by sampling a vector $s$ of length $N$ generally from a ternary distribution with Hamming weight $h$.
+The public key is constructed by sampling $a_{0}$ from $\mathbb{Z}_{Q}$ (i.e., from the set of positive integers less than the modulus $Q$) and an error term $e_{0}$ from a discrete Gaussian distribution with standard deviation $\sigma=3.2$, as follows:
 $pk=([-a_{0}s+e_{0}]_{Q},a_{0})=(p_{0},p_{1})$
-where the notation $[\cdots]_{Q}$ denotes reduction modulo $Q$.
+where the notation $\cdots]_{Q}$ denotes reduction modulo $Q$.
 Encoding
@@
 Different libraries implementing CKKS vary in how they perform this transformation using IFFT/FFT-based methods.
-In all cases, the standard cyclic transform must be adapted to a negacyclic version to ensure operations remain within the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$ used by the scheme.
-The transforms employed by HEAAN 1.0 and OpenFHE 1.2.3 are negacyclic and include scheme-specific optimizations, commonly referred to as specialized FFT/IFFT.
+In all cases, the standard cyclic transform must be adapted to a negacyclic version to ensure that operations remain within the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$ used by the scheme.
+The transforms employed by the most adopted CKKS implentations are negacyclic and include scheme-specific optimizations, commonly referred to as specialized FFT/IFFT.
 These versions incorporate several improvements originally proposed by Bernstein [15, 16].
 One of their main advantages is that they allow the transform to be applied directly to vectors of size $\leq N/2$, without expanding them to size $N$, which significantly reduces computational cost.
-Given a real-valued input vector of size $n\leq N/2$, the encoding process begins by padding the vector with zeros to reach the next power of two, denoted $n^{\prime}$.
+Given a complex-valued input vector of size $n\leq N/2$, the encoding process begins by padding the vector with zeros to reach the next power of two, denoted $n^{\prime}$.
 A specialized inverse FFT is then applied, producing a complex vector of length $n^{\prime}$.
-From this, a polynomial of degree $N-1$ is constructed.
-First, the gap parameters $=\frac{N/2}{n^{\prime}}$ is computed, and the lower half of the polynomial (coefficients of degree 0 to $N/2-1$) is filled with the real parts of the transformed vector, with assignments spaced at intervals of size gap.
+From this, a polynomial of degree N1N-1 is constructed.
+First, the gap parameter $gap=\frac{N/2}{n^{\prime}}$ is computed, and the lower half of the polynomial (coefficients of degree 0 to N/21N/2-1) is filled with the real parts of the transformed vector, with assignments spaced at intervals of size gap.
+The upper half does the same using the imaginary part of the transformed vector.
 Unassigned coefficients are set to zero.
-Finally, as is illustrated in Fig. 3, each coefficient is multiplied by $\Delta$, and then rounded to the nearest integer using the coordinate-wise randomized rounding method described in [21].
-Figure 3: Encoding scheme in CKKS using HEAAN/OpenFHE.
-The example shows an input vector with $n=2$ values, represented in a polynomial ring of size $N=8$, where the gap between consecutive coefficients is 2.
+This process is depicted in Figure 3.
+As part of this encoding stage, each coefficient is multiplied by $\Delta$, and then rounded to the nearest integer using the coordinate-wise randomized rounding method described in [17].
+Figure 3: Standard encoding scheme in CKKS. The example shows an input vector with $n=2$ values, represented in a polynomial ring of size $N=8$, where the gap between consecutive coefficients is 2.
 Encryption
@@
 $c_{sk}=([m+a\times s+e]_{Q},[-a]_{Q})=(c_{0},c_{1})$ (2)
-where $a$ is a polynomial sampled from a uniform distribution, $e$ is a noise term sampled from a discrete Gaussian distribution, and the notation $[\cdots]_{Q}$ denotes reduction modulo $Q$.
-An alternative way is encrypting by using the public key, $pk=(p_{0},p_{1})$. In this case, encryption is defined as:
+where $a$ is a polynomial sampled from a uniform distribution, $e$ is a noise term sampled from a discrete Gaussian distribution, and the notation $\cdots]_{Q}$ denotes reduction modulo $Q$.
+An alternative procedure is to encrypt by using the public key, $pk=(p_{0},p_{1})$. In this case, encryption is defined as:
 $c_{pk}=([m+p_{0}\times v+e_{1}]_{Q},[p_{1}\times v+e_{2}]_{Q})=(c_{0},c_{1})$ (3)
@@
 Decryption
-Decryption in both cases (secret key and public key encryption) uses the secret key $sk=(1,s)$ and the encrypted text $(c_{0},c_{1})$ as follows:
-$m^{\prime}=[c_{0}+c_{1}\times s]_{Q}$ (4)
+Decryption in both secret key and public key cases uses the secret key $sk=(1,s)$ and the encrypted text $(c_{0},c_{1})$ as follows:
+$m=[c_{0}+c_{1}\times s]_{Q}$ (4)
 Decoding
-To decode a plaintext polynomial $m$ with $N$ coefficients, one must reconstruct a complex-valued vector of $n^{\prime}$ elements by reversing the encoding process illustrated in Fig.
-3.
+Decoding a plaintext polynomial $m$ with $N$ coefficients implies reconstructing a complex-valued vector of $n^{\prime}$ elements by reversing the encoding process illustrated in Figure 3.
 The $i$-th element ($i\leq N/2$) of the resulting vector is formed by using the real part of the $i$-th coefficient of the plaintext, and the imaginary part from the coefficient at position $i+N/2$.
 This complex vector is then divided by $\Delta$, and a specialized FFT is applied to recover the approximate original input.
-Due to rounding effects during encoding and decoding, the final result may contain small approximation errors.
+Even when simply encoding and decoding an input vector without performing any computation, small approximation errors may arise due to rounding effects.
 Corollary 3 in [18] shows that the worst-case distance between the input vector $z$ and the output vector $z^{\prime}$ (after encoding and decoding) at each position is approximately: $N/(\pi\times\Delta)$.
-[…]
-Security Against Secret Key Attacks
-To defend against secret key recovery attacks, OpenFHE performs additional operations before applying the specialized FFT during the decoding process.
-Specifically, a Gaussian error is injected prior to this step, with its standard deviation determined by the vector being decoded.
-This mechanism is critical: in our bit-flip experiments–when security analysis is not disabled–the software throws an exception during this stage, resulting in premature termination of execution.
-This behavior demonstrates the role of these protective measures in preventing potential side-channel or fault-injection attacks that could compromise the secret key.
-II-D RNS and NTT Optimizations
-In the full RNS/NTT variant of CKKS [3], the Residue Number System (RNS) is used to split large polynomial coefficients into multiple smaller values, enabling more efficient computations.
-This approach allows arithmetic in $\mathbb{Z}_{Q}$ to be performed using standard machine word sizes, by working with small moduli that simplify computation.
-The magnitude of the modulus $Q$ can range from hundreds to thousands of bits, depending on the use case.
-However, modern computing systems typically operate with word sizes of 64 bits or less, making direct operations on such large coefficients impractical.
-To address this, polynomials (keys, plaintexts, ciphertexts) are represented using an equivalent decomposition in the RNS form.
-Each coefficient is represented modulo a small prime $q_{i}$, to generate $Q$ like Eq.
-1
-This decomposition enables addition and multiplication in $\mathbb{Z}_{Q}$ using standard machine word operations.
-All other parameters, as well as the encoding and decoding procedures, remain consistent with the vanilla CKKS variant.
-Polynomial multiplications are performed efficiently using the Number Theoretic Transform (NTT), which reduces the computational complexity of the operation from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$.
-Addition of polynomials and multiplication by a scalar are both $\mathcal{O}(n)$ operations in both the coefficient and evaluation (NTT) representations.
-By default, the CKKS scheme assumes that polynomials are represented in the NTT domain (evaluation representation).
-When certain operations require polynomials in coefficient form, an inverse NTT (iNTT) is applied to transform ciphertexts from the evaluation domain back to the coefficient domain.
-However, within the scope of this work, as all computations are performed on the client side, this change of representation does not occur.
-The procedures for encoding, encryption, decryption, and decoding remain virtually unchanged, except that all operations are now performed using the RNS and NTT representations.
+With a sufficiently large $\Delta$, this difference becomes negligible.
+Given that a typical ring size $N$ ranges from $2^{12}-2^{16}$, using a scaling factor on the order of $2^{20}$ is enough to ensure a negligible approximation error.
 III Error Resilience Analysis
-This section presents a theoretical analysis of the error induced by a single bit-flip at different stages of the CKKS scheme.
-We address both the vanilla and the full RNS variant versions.
-Our aim is:
-  • Deriving analytical bounds in the case without RNS or NTT.
-  • Identifying the most sensitive coefficients.
-  • Providing an intuitive description of error behavior in practical implementations.
-Although a closed-form analytical bound can not be derived for the full RNS variant version due to the complexity of modular representations, a qualitative analysis of the observed patterns is provided.
-III-A Error Model for a Single Bit-Flip
-We adopt a single bit-flip model, where a single bit is altered. In practice, such faults may be caused by electromagnetic interference, hardware failures, or cosmic rays [12, 15, 23, 13, 24, 17].
+This section presents a theoretical analysis of the error induced by a single bit flip at different stages of the CKKS scheme, in its vanilla and full RNS variants. The goal of this analysis is:
+  • Determining analytical error bounds in the vanilla variant (no RNS/NTT).
+  • Identifying the most sensitive polynomial coefficients.
+  • Enabling a better understanding of error behavior in practical implementations.
+Although a closed-form analytical bound can not be derived for the full RNS variant due to the complexity of modular representations, a qualitative analysis of the observed patterns is provided.
+III-A Error Model for a Single Bit Flip
+We adopt a single bit-flip (soft error) model, where a single bit is altered.
+In practice, such faults may be caused by electromagnetic interference [7, 10], hardware failures [19, 8], or cosmic rays [20, 12].
 Let $p$ be the original value of a coefficient, either in a plaintext polynomial $\boldsymbol{m}$ or in one of the ciphertext polynomials $(c_{0},c_{1})$.
@@
 $e_{j}=p_{j}^{\prime}-p$ (5)
-We model this fault as an error vector $\boldsymbol{e}_{i,j}\in\mathbb{R}^{N}$ whose only non-zero component corresponds to bit $j$ of coefficient $i$:
-$\mathbf{e}_{i,j}=(0,\dots,0,e_{j},0,\dots,0)$ (6)
-This model enables us to track how a single bit-flip propagates through the cryptographic pipeline.
+We model this fault as an error polynomial $\boldsymbol{e}_{i,j}\in\mathbb{Z}[X]$, whose only non-zero coefficient is the (i1)(i-1)-th, corresponding to an error at bit $j$ of coefficient i1i-1:
+$\boldsymbol{e}_{i,j}=e_{j}\,X^{i-1}$ (6)
+This model enables us to track how a single bit flip propagates through the cryptographic pipeline.
 III-B CKKS Encoding via DFT
-As discussed in Section II, one of the key components of encoding and decoding in CKKS is FFT.
+As discussed in Section II, the Fast Fourier Transform (FFT) is one of the key components of encoding and decoding in CKKS.
 Although CKKS uses a specialized inverse FFT for computational efficiency, this transformation is typically implemented using optimized algorithms that obscure its underlying linear structure.
-Instead, we rely on the negacyclic Discrete Fourier Transform (DFT)–an algebraically equivalent transformation explicitly expressed as matrix multiplication. This formulation allows us to:
+Instead, we rely on the negacyclic Discrete Fourier Transform (DFT), an algebraically equivalent transformation explicitly expressed as matrix multiplication.
+This formulation allows us to:
   1. 1.
@@
     Derive analytical bounds on the magnitude of the error without dealing with the internal optimizations of FFT implementations.
-While this choice sacrifices performance, it offers significant pedagogical advantages, allowing for an intuitive understanding of how bit-flip errors are amplified or distributed during encoding.
+While this choice sacrifices performance, it offers significant pedagogical advantages, allowing for an intuitive understanding of how bit-flip errors are amplified or distributed during the encoding stage.
 III-B1 Encoding via Negacyclic DFT
@@
 $W^{-1}=\frac{1}{N}W^{\dagger}$
-In order to use the DFT and its inverse for encoding to obtain results equivalent to those produced by implementations such as HEAAN and OpenFHE, some preprocessing is required: the input vector must be resized to length $N$ and reordered to satisfy the requirements of those systems.
-However, these adjustments do not affect the behavior of the error model under analysis and will be omitted for simplicity.
+To use the DFT and its inverse for encoding to obtain results equivalent to those produced by standard implementations, some preprocessing is required: the input vector must be resized to length $N$ and reordered.
+However, these adjustments do not affect the behavior of the error model under analysis and are omitted for simplicity.
 In general, for an input vector of $n\leq N/2$ elements, it is first zero-padded to length $N/2$, and then mirrored with its complex conjugate to reach a total length of $N$.
@@
 Finally, the output undergoes the same reordering process as in encoding, and the $n$ components of $z^{\prime}$ are extracted.
-Since the DFT output is complex-valued while OpenFHE supports only real-valued encryption, the real part is extracted to obtain the final approximation of z
-III-C Theoretical Estimation of Encoding Error in Vanilla CKKS
+Since the DFT output is complex-valued, the real part is extracted to obtain the final approximation of z
+III-C Theoretical Estimation of Encoding Error on vanilla CKKS
 This section focuses on the impact of single-bit errors during the encoding phase of the CKKS scheme.
-Let us consider the case in which an error $\boldsymbol{e}_{i,j}$ is introduced by flipping a single bit in coefficient $i$ of the plaintext vector $\boldsymbol{m}$, specifically at the $j$-th bit position.
-The resulting perturbed plaintext is $\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j}$, and decoding proceeds as in Eq.
-9:
+Let us consider the case in which an error $\boldsymbol{e}_{i,j}$ is introduced by flipping a single bit in the coefficient $i$ of the plaintext vector $\boldsymbol{m}$, specifically at the $j$-th bit position.
+The resulting perturbed plaintext is $\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j}$, and decoding proceeds as in Equation 9:
 $\begin{split}\mathbf{z^{\prime\prime}}&=W\times\frac{\mathbf{m^{\prime}}}{\Delta}=W\times\frac{\mathbf{m}+\mathbf{e}_{i,j}}{\Delta}\\ &=W\times\frac{\mathbf{m}}{\Delta}+W\times\frac{\mathbf{e}_{i,j}}{\Delta}\\ &=\mathbf{z^{\prime}}+W\times\frac{\mathbf{e}_{i,j}}{\Delta}\end{split}$
-Since a single bit-flip affects only one coefficient, the resulting error is localized.
+Since a single bit flip affects only one coefficient, the resulting error is localized.
 The difference between the decoded vectors $\boldsymbol{z^{\prime}}$ and $\boldsymbol{z^{\prime\prime}}$ is given by:
@@
 III-C1 Bit-Flip Injection during Encoding
-To better understand the analytical expression of the error in Eq.
-(10), we simulate flipping each individual bit in the plaintext coefficients.
+To better understand the analytical expression of the error in Equation 10, we simulate flipping each bit in the plaintext coefficients.
 Let $j\in\{0,\dots,63\}$ denotes the bit position in a 64-bit unsigned integer.
 For each iteration, we flip the $j$-th bit of a coefficient, introducing an error with integer magnitude $e_{j}=2^{j}$.
-This represents the maximum jump in value induced by flipping bit $j$ in a coefficient, which is stored as an unsigned integer–as in the case of plaintext coefficients.
-We repeat this process for all coefficients, grouping the results into $N=4$ blocks of 64 iterations each, and use $\Delta=2^{50}$.
-Let $\boldsymbol{z}^{\prime}$ be the decoded vector without errors, and $\boldsymbol{z}^{\prime\prime}$ the result after applying the bit-flip and decoding.
-Using Eq.
-(10), we define:
+This represents the maximum jump in value induced by flipping bit $j$ in a coefficient, which is stored as an unsigned integer (which is the case of plaintext coefficients).
+We repeat this process for all coefficients, grouping the results into $N=4$ blocks of 64 iterations each, and using $\Delta=2^{50}$.
+Let $\boldsymbol{z}^{\prime}$ be the decoded vector without errors, and $\boldsymbol{z}^{\prime\prime}$ the result after applying the bit flip and decoding.
+Using Equation 10, we express the error using the $L_{2}$ norm:
 $L_{2}(i,j)=\bigl{\lVert}Real(W\times\boldsymbol{e}_{i,j}/\Delta)\bigr{\rVert}_{2}$ (11)
 where $\mathbf{e}_{i,j}$ is a zero vector of length $N$ with value $e_{j}$ at position $i$.
-We also take the real part of the vector before computing the L2 norm, to simulate the encoding of real-valued data, as in OpenFHE, in contrast to HEAAN, which allows complex inputs.
+We also take the real part of the vector before computing the $L_{2}$ norm, to simulate the encoding of real-valued data.
 This step is important not only to maintain consistency with the backend used in later sections, but also because it exposes a specific behavior that would otherwise not manifest when keeping the complex part.
-Figure 4: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding.
-The horizontal axis shows the bit-flip position $i*64+j$, and the vertical axis shows $L_{2}(i,j)$.
-Ring dimension $N=4$; scaling factor $\Delta=2^{50}$.
+Figure 4: $L_{2}$-norm of the error, measured as the difference between the golden (no bit-flip) output and the output obtained with a single bit-flip in the plaintext during encoding.
+The horizontal axis shows the bit-flip position.
+Parameters: $N=4$ and $\Delta=2^{50}$.
 Figure 4 shows the values of $L_{2}(i,j)$ as a function of the bit-flip position.
@@
 For each case, we plot $L_{2}(i,j)$ using a semi-logarithmic scale on the $y$-axis for easier comparison.
-Figure 5: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding, for various scaling factors $\Delta$.
-The horizontal axis indicates the bit position $j$; the vertical axis is in semi-log scale.
+Figure 5: $L_{2}$-norm of the error after a single bit-flip in the plaintext during encoding, for various scaling factors $\Delta$.
+The horizontal axis indicates the bit position $j$; the vertical axis is on a semi-log scale.
 Ring dimension: $N=4$.
-As shown in Fig.
-5, increasing $\Delta$ significantly reduces the error magnitude and increases the number of resilient bits.
+As shown in Figure 5, increasing $\Delta$ significantly reduces the error magnitude and increases the number of resilient bits.
 This occurs because the impact of errors becomes noticeable only when the flipped bit corresponds to or exceeds the magnitude of $\Delta$.
 In other words, the larger the scaling factor, the more initial bits remain unaffected, yielding smaller error norms and higher bit-level fault tolerance.
-This empirical observation is analytically confirmed by equation 11, which shows that the error is inversely proportional to the scaling factor.
+This empirical observation is confirmed analytically by equation 11, which shows that the error is inversely proportional to the scaling factor.
 When an error $e_{i,j}$ is divided by $\Delta$, noise components with magnitudes smaller than the bit precision of $\Delta$ are effectively suppressed in the final result.
@@
 However, this resilience comes at the cost of reduced fractional precision in the encoded message (see Section II).
-The selection of $\Delta$ therefore requires careful consideration of the trade-off between error tolerance and numerical precision, balancing the competing demands of fault resilience and computational accuracy.
+The selection of $\Delta$ therefore requires careful consideration, balancing the competing demands of fault resilience and computational accuracy.
 III-D Theoretical Estimation of Error during Encryption using Vanilla CKKS
-This section analyzes three scenarios involving a single bit-flip in vanilla CKKS:
-  1. 1.
-    A bit-flip in the plaintext (before encryption).
-  2. 2.
-    A bit-flip in the ciphertext (in one of the ciphertext polynomials).
-  3. 3.
-    A bit-flip in the plaintext (before decoding).
-Under public-key encryption, a plaintext vector $\mathbf{m}$ is encoded into a polynomial $\boldsymbol{m}(X)$, encrypted as in Eq. 3, and decrypted using Eq. 4.
-III-D1 Bit‐Flip in the Plaintext Before Encryption
-Suppose a bit-flip introduces an error $\boldsymbol{e}_{i,j}$ in the $i$-th coefficient at bit position $j$, so the plaintext becomes:
-$\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j},\quad\boldsymbol{e}_{i,j}=e_{j}\,X^{i-1}$ (12)
-Encrypting $\boldsymbol{m}^{\prime}$ using the public key (as in Eq. 3) yields:
-$\begin{split}ct^{\prime}&=\bigl{(}[\,\boldsymbol{m}^{\prime}+p_{0}v+e_{1}\,]_{Q},\;[\,p_{1}v+e_{2}\,]_{Q}\bigr{)}\\ &=\bigl{(}[\,\boldsymbol{m}+e_{j}\,X^{i-1}+p_{0}v+e_{1}\,]_{Q},\;c_{1}\bigr{)}=\bigl{(}c_{0}^{\prime},\,c_{1}\bigr{)}.\end{split}$ (13)
-The error affects only the ciphertext component $c_{0}^{\prime}$ in coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic. Upon decryption (using Eq. 4):
-$\boldsymbol{m}^{\prime\prime}=[\,c_{0}^{\prime}+c_{1}\,s\,]_{Q}=\boldsymbol{m}+e_{j}\,X^{i-1}+(e_{1}+e_{2}\,s)$ (14)
+Under public-key encryption, a vector $\mathbf{m}$ is encoded into a polynomial $\boldsymbol{m}(X)$, encrypted as in Equation 3, and decrypted using Equation 4.
+III-D1 Bit Flip in the Plaintext Before Encryption
+Assuming that a bit flip introduces an error $\boldsymbol{e}_{i,j}$, as previously explained, the plaintext becomes:
+$\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j}$ (12)
+Encrypting $\boldsymbol{m}^{\prime}$ using the public key (Equation 3) yields:
+$\begin{split}ct^{\prime}&=\bigl{(}[\,\boldsymbol{m}^{\prime}+p_{0}v+e_{1}\,]_{Q},\;[\,p_{1}v+e_{2}\,]_{Q}\bigr{)}\\ &=\bigl{(}[\,\boldsymbol{m}+\boldsymbol{e}_{i-1,j}+p_{0}v+e_{1}\,]_{Q},\;c_{1}\bigr{)}=\bigl{(}c_{0}^{\prime},\,c_{1}\bigr{)}.\end{split}$ (13)
+The error affects only the ciphertext component $c_{0}^{\prime}$ in the coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic.
+Decryption (Equation 4) of such erroneous ciphertext yields:
+$\boldsymbol{m}^{\prime\prime}=[\,c_{0}^{\prime}+c_{1}\,s\,]_{Q}=\boldsymbol{m}+\boldsymbol{e}_{i-1,j}+(e_{1}+e_{2}\,s)$ (14)
 However, the error magnitude grows exponentially with the flipped bit position, as each bit represents a successive power of 2 in the binary representation.
-Consequently, an exhaustive bit-flip analysis across all coefficient positions, with L2 norm computation for each perturbation, yields the exponential error scaling pattern illustrated in Fig.
-4 (SectionIII-C).
-that the bit-flip impact remains localized to the position $i$ and scales linearly with $e_{j}$.
-The error induced by the bit-flip affects only the ciphertext component $c_{0}^{\prime}$ in coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic.
-Decoding this plaintext will therefore produce an error consistent with that discussed in Section III-C.
-III-D2 Bit‐Flip in the Ciphertext
-If the bit-flip occurs in the ciphertext, two sub-cases arise:
+Consequently, an exhaustive bit-flip analysis across all coefficient positions, with L2 norm computation for each perturbation, yields the exponential error scaling pattern illustrated in Figure 4 (SectionIII-C).
+The bit-flip impact remains localized to the position $i$ and scales linearly with $e_{j}$.
+III-D2 Bit Flip in the Ciphertext
+If the bit flip occurs in the ciphertext, two sub-cases arise:
 Flip in $c_{0}$
-For a perturbed ciphertext $ct^{\prime}=(c_{0}+e_{j}\,X^{i-1},\;c_{1})$, the behavior mirrors the previous plaintext case: the error remains localized in coefficient $i$ and results in a decoding error proportional to $e_{j}$.
+For a perturbed ciphertext $ct^{\prime}=(c_{0}+\boldsymbol{e}_{i-1,j},\;c_{1})$, the behavior mirrors the previous plaintext case: the error remains localized in coefficient $i$ and results in a decoding error proportional to $e_{j}$.
 Flip in $c_{1}$
-For a perturbed ciphertext $ct^{\prime}=(c_{0},\;c_{1}+e_{j}\,X^{i-1})$, decryption proceeds as:
-$\begin{split}\boldsymbol{m}^{\prime\prime}&=[\,c_{0}+(c_{1}+e_{j}\,X^{i-1})\,s\,]_{Q}\\ &=[\,(c_{0}+c_{1}\,s)+(e_{j}\,X^{i-1}\,s)\,]_{Q}\\ &=[\,(\boldsymbol{m}+e_{1}+e_{2}\,s)+(e_{j}\,X^{i-1}\,s)\,]_{Q}\end{split}$
-where $e_{j}X^{i-1}\,s$ involves polynomial multiplication.
+For a perturbed ciphertext $ct^{\prime}=(c_{0},\;c_{1}+\boldsymbol{e}_{i-1,j})$, decryption proceeds as:
+$\begin{split}\boldsymbol{m}^{\prime\prime}&=[\,c_{0}+(c_{1}+\boldsymbol{e}_{i-1,j})\,s\,]_{Q}\\ &=[\,(c_{0}+c_{1}\,s)+(\boldsymbol{e}_{i-1,j}\,s)\,]_{Q}\\ &=[\,(\boldsymbol{m}+e_{1}+e_{2}\,s)+(\boldsymbol{e}_{i-1,j}\,s)\,]_{Q}\end{split}$
+where $\boldsymbol{e}_{i-1,j}\,s$ involves polynomial multiplication.
 Thus, the error spreads across all $N$ coefficients.
@@
 The maximum magnitude is bounded by $|e_{j}|\cdot\|s\|_{\infty}$.
-III-D3 Bit‐Flip in the Plaintext Before Decoding
-If a bit-flip is introduced in coefficient $i$ of a plaintext $\boldsymbol{m}$, after decryption but before decoding, the effect is analogous to the case analyzed in Section III-C.
-This follows directly since the plaintext being decoded is equivalent to:
-$\boldsymbol{m}^{\prime}=\boldsymbol{m}+e_{j}\,X^{i-1}$ (15)
+III-D3 Bit flip in the Plaintext Before Decoding
+If a bit flip is introduced in the coefficient $i$ of a plaintext $\boldsymbol{m}$, after decryption but before decoding, the effect is analogous to the case analyzed in Section III-C, since the plaintext being decoded is equivalent to:
+$\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i-1,j}$ (15)
 which is the same structure previously analyzed.
@@
 Each coefficient $p\in\mathbb{Z}_{Q}$ is represented by its residues:
-$r_{k}=p\bmod q_{k}\quad\text{for }k=0,\dots,L$
-When applied to all coefficients of a polynomial (from an encoded message or ciphertext), this yields $L$ residue polynomials, commonly referred to as limbs.
-Reconstruction via the Chinese Remainder Theorem (CRT), as shown in Mohan [25, Chapter 5], is:
+$r_{k}=p\bmod q_{k}\quad\text{for }k=0,\dots,L-1$
+When applied to all coefficients of a polynomial (either in the plaintext or in the ciphertext), this yields $L$ residue polynomials, commonly referred to as limbs.
+Reconstruction via the Chinese Remainder Theorem (CRT), as shown in Mohan [21, Chapter 5], is:
 $p=\left(\sum_{k=1}^{L}r_{k}\left[\left(\frac{1}{Q_{k}}\right)\bmod q_{k}\right]Q_{k}\right)\text{ mod }Q$ (16)
-where
-$Q_{k}=Q/q_{k}$
-and $\frac{1}{Q_{k}}$ are the multiplicative inverse of $Q_{k}$.
-A bit-flip on the $j$ bit affecting the $k$-th residue changes it to $r_{k}^{\prime}=r_{k}+e_{k,j}$, which introduces an error into the reconstructed value:
-$p^{\prime}=p+e_{k,j}\,Q_{k}\,\left[\left(\frac{1}{Q_{k}}\right)\bmod q_{k}\right]\pmod{Q}$ (17)
-This error can be as large as $Q_{k}$, potentially spanning hundreds or thousands of bits, even though the bit-flip affected only one bit in a single residue.
+where $Q_{k}=Q/q_{k}$ and $\frac{1}{Q_{k}}$ are the multiplicative inverse of $Q_{k}$.
+A bit flip on the $j^{th}$ bit affecting the $k$-th residue changes it to $r_{k}^{\prime}=r_{k}+e_{k,j}$, which introduces an error into the reconstructed value:
+$p^{\prime}-p=e_{k,j}\,Q_{k}\,\left[\left(\frac{1}{Q_{k}}\right)\bmod q_{k}\right]\pmod{Q}$ (17)
+This error can be as large as $Q_{k}$, potentially spanning hundreds or thousands of bits, even though the bit flip originally affected a single bit in a single residue.
 III-F Theoretical Estimation of Error in NTT
-[…]
+To accelerate encoding and decoding operations, CKKS uses the negacyclic Number Theoretic Transform (NTT) over the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$.
+Let $\psi\in\mathbb{Z}_{q}$ be a primitive $2N$-th root of unity satisfying:
 $\psi^{2}\equiv\xi,\quad\psi^{N}\equiv-1\pmod{q}$,
-where $\xi$ is a primitive $N$-th root of unity. The NTT matrix is then defined as:
+where $\xi$ is a primitive $N$-th root of unity. The NTT matrix is defined as:
 $W_{NTT}=\begin{pmatrix}\psi^{2(0\times 0).0}&\psi^{2(0\times 1).1}&\cdots&\psi^{2(0\times N-1).N-1}\\ \psi^{2(1\times 0).0}&\psi^{2(1\times 1).1}&\cdots&\psi^{2(1\times N-1).N-1}\\ \psi^{2(2\times 0).0}&\psi^{2(2\times 1).1}&\cdots&\psi^{2(2\times N-1).N-1}\\ \vdots&\vdots&\ddots&\vdots\\ \psi^{2(N-1\times 0).0}&\psi^{2(N-1\times 1).1}&\cdots&\psi^{2(N-1\times N-1).N-1}\end{pmatrix}$
-We denote its inverse as $W^{-1}_{NTT}$. When a single bit-flip occurs at bit position $j$ within coefficient $i$, the induced perturbation after NTT transformation becomes:
-$\mathbf{m}^{\prime}-\mathbf{m}=W_{NTT}\,\mathbf{e}_{i,j}.$
+We denote its inverse as $W^{-1}_{NTT}$.
+When a single bit flip occurs at bit position $j^{th}$ within coefficient $i$, the induced perturbation after NTT transformation becomes $\mathbf{m}^{\prime}-\mathbf{m}=W_{NTT}\,\mathbf{e}_{i,j}$.
 Therefore, the $L_{2}$ norm of the error is:
 $\lVert\mathbf{m}^{\prime}-\mathbf{m}\rVert_{2}=\lVert W_{NTT}\,\mathbf{e}_{i,j}\rVert_{2}.$
+This behavior parallels that of the Discrete Fourier Transform (DFT), except that the NTT operates over $\mathbb{Z}_{q}$ with moduli $q$ typically of 306030-60 bits.
+The twiddle factors, being roots of unity modulo $q$, can amplify errors during the transformation, thereby increasing the error.
+Figure 6 shows the values of $L_{2}(i,j)$ for $N=4$ and 64-bit coefficients in the NTT case.
+The horizontal axis indicates the bit-flip position $i\times 64+j$, and the vertical axis shows $L_{2}(i,j)$ on a logarithmic scale.
+Figure 6: $L_{2}$-norm of the error after a single bit-flip in the plaintext during encoding.
+The horizontal axis shows the modified bit position $i\times 64+j$; the vertical axis shows the $L_{2}(i,j)$ norm on a logarithmic scale.
+Ring dimension: $N=4$.
+As Figure 6 shows, even flips in the least significant bits cause the norm (error) to exceed four orders of magnitude–producing errors large enough to make correct decoding practically impossible, as will be further illustrated in Section V.
+III-G Summary
+Bit flip in the plaintext before encryption
+the error remains localized in a single coefficient (the one where the bit flip occurred), with a magnitude proportional to $e_{j}=2^{j}$, depending on the position $j$ of the flipped bit.
+Bit flip in the ciphertext ($c_{0}$)
+identical to the plaintext case–the error remains localized after decryption, with magnitude proportional to the bit position $j$.
+Bit flip in the ciphertext ($c_{1}$)
+the error is multiplied by the secret key $s$, becoming distributed across all $N$ coefficients; its maximum magnitude is $\mathcal{O}(|e_{j}|\cdot\|s\|_{\infty})$.
+Bit flip in the plaintext before decoding
+behaves like a localized error in a single coefficient of the encoded polynomial–effectively equivalent to the encoding case, plus minor encryption noise.
+RNS
+a bit flip in a single limb (residue) can result in a reconstruction error of order $\mathcal{O}(Q_{k})$, where $Q_{k}=Q/q_{k}$ is typically a large value.
+Negacyclic NTT
 […]
-Figure 6 shows the values of $L_{1}(i,j)$ for $N=4$ and 64-bit coefficients in the NTT case.
-The horizontal axis indicates the bit-flip position $i\cdot 64+j$, and the vertical axis shows $L_{2}(i,j)$ on a logarithmic scale.
-Figure 6: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding.
-The horizontal axis shows the modified bit position $i*64+j$; the vertical axis shows the $L_{2}(i,j)$ norm on a logarithmic scale.
-Ring dimension: $N=4$.
-As the Fig.
-6 shows, even the flips in the least significant bits cause the norm to exceed four orders of magnitude—producing errors large enough to make correct decoding practically impossible, as will be further illustrated in Section V.
-III-G Summary
-The following points summarize the theoretical impact of a single bit-flip at different stages of the CKKS encryption pipeline, both in its vanilla and optimized implementations:
-  • Bit-flip in the plaintext before encryption: the error remains localized in a single coefficient where the bit-flip occurred, with a magnitude proportional to $e_{j}=2^{j}$, depending on the flipped bit’s position $j$.
-  • Bit-flip in the ciphertext ($c_{0}$): identical to the plaintext case—the error remains localized after decryption, with magnitude again proportional to the bit position $j$.
-  • Bit-flip in the ciphertext ($c_{1}$): the error is multiplied by the secret key $s$, becoming distributed across all $N$ coefficients; its maximum magnitude is $\mathcal{O}(e_{j}\|s\|_{\infty})$.
-  • Bit-flip in the plaintext before decoding: behaves like a localized error in a single coefficient of the encoded polynomial–effectively equivalent to the encoding case, plus minor encryption noise.
-  • RNS: a bit-flip in a single limb (residue) can result in a reconstruction error of order $\mathcal{O}(Q_{k})$, where $Q_{k}=Q/q_{k}$ is typically a large value.
-[…]
-IV Experimental Design
-This section describes the experimental design used to quantify the resilience of the CKKS homomorphic encryption scheme under single bit-flip fault injections.
-IV-A Test Environment
-[…]
-  • HEAAN 1.0: a vanilla implementation without RNS or NTT, using arbitrary-precision coefficients managed by the NTL library (a C++ library for number theory and arbitrary-precision arithmetic).
-  • OpenFHE 1.2.3: a modular implementation with native support for RNS and NTT, using 64-bit internal coefficients (uint64).
-Although more recent versions of HEAAN include RNS and NTT optimizations, version 1.0 allows a cleaner comparison to a vanilla CKKS.
-Conversely, OpenFHE represents the optimized variant, due to its widespread use in the field of homomorphic encryption [26] and its active support for high-performance execution environments.
-Minor modifications were made to both libraries to gain full control over the selection of pseudo-random number generator (PRNG) seeds.
-This ensured strict reproducibility of experiments and allowed for precise isolation of the effects caused by each individual bit-flip.
-In the case of HEAAN 1.0, we have these additional implementation details:
-  • The bit-length of each coefficient is not fixed in advance; it depends on internal combinations of $q_{0}$ and $\Delta$, and may vary depending on the PRNG seed used by the scheme.
-  • Due to NTL implementation, bits larger than $Q$ are ignored in internal operations, which limits the scope of bit-level analysis.
-OpenFHE has the following implementation details:
-  • Uses 64-bit integer coefficients (uint64_t) without arbitrary-precision arithmetic.
-  • The initial modulus $q_{0}$ can be explicitly configured by the user.
+IV Methodology
+This section describes the methodology used to experimentally quantify the resilience of the CKKS scheme under single bit-flip fault injections.
+IV-A Scheme Definition and Operational Assumptions
+We consider four experimental configurations using CKKS:
+  • Vanilla mode: All low-level performance optimizations are disabled; arithmetic occurs in a single large modulus in coefficient space (no RNS–single limb–and no NTT domain representation).
+  • RNS-only mode: RNS decomposition is enabled (multi-limb representation), while NTT is bypassed.
+  • NTT-only mode: NTT is enabled to accelerate polynomial multiplication / convolution, while RNS decomposition is disabled (single-limb modulus).
+  • RNS+NTT mode: Fully optimized configuration (default in OpenFHE) with both RNS decomposition and NTT transform.
+This setup allows us to isolate the individual and combined effects of RNS and NTT optimizations on fault propagation and resilience, while maintaining compatibility with standard security parameters.
+In preliminary tests, we observed that both the NTT-only and the RNS+NTT configurations exhibit similar error amplification even for single-bit faults, in line with our theoretical analysis (Section III).
+In other words, a single bit flip at any stage of these two cases leads to undecryptable results.
+As a result, we focus our analysis on the other two cases (vanilla and RNS-only modes), where more interesting behavior can be observed and meaningfully characterized.
+IV-B Fault / Error Injection Model
+As in our theoretical analysis (Section III-A), we assume a transient single-bit fault model, motivated by soft errors such as radiation-induced bit flips in processor registers or memory buffers.
+Specifically, we inject a single bit flip in one coefficient of the output of a given operation during each execution.
+In the case of Encode and Decode, we inject the fault into the plaintext representation; for Encrypt and Decrypt, we inject it into the ciphertext.
+This approach models a fault affecting the destination register or output memory of these functions, without modifying the inputs or intermediate computations.
+Figure 7: An input is first encoded and then encrypted.
+A single bit flip is applied to one coefficient at a time and at one bit position.
+The ciphertext is then decrypted and decoded, and the resulting output is compared against the unmodified reference output using the $L_{2}$-norm.
+For illustration, the ciphertext is represented as a single polynomial.
+Figure 7 illustrates how the fault is injected at a single stage per experiment, directly on the polynomial representation (plaintext or ciphertext), simulating a post-operation corruption that propagates through subsequent stages of the computation.
+IV-C Implementations and Instrumentation
+IV-C1 CKKS Libraries
+We use two different CKKS implementations: HEAAN 1.0.
+First publicly available CKKS implementation.
+This specific version does not support RNS and NTT, which makes it suitable for the evaluation of the vanilla mode.
+It uses arbitrary-precision coefficients managed by the NTL library.
+OpenFHE 1.3.0: highly-optimized open-source (de facto) standard implementation of CKKS [22].
+It natively supports RNS and NTT, using 64-bit internal coefficients.
+We minimally adapt both libraries to gain complete control over the selection of pseudo-random number generator (PRNG) seeds.
+This ensures strict reproducibility of experiments and allows precise isolation of the effects caused by individual bit flips.
+In the case of HEAAN 1.0, we consider the following implementation details:
+  • The bit length of each plaintext coefficient is not fixed in advance; it depends on internal combinations of $q_{0}$ and $\Delta$, and may vary depending on the PRNG seed used by the scheme.
+  • The ciphertext bit length is of length $q_{0}$.
+  • Due to NTL implementation constraints, bit positions larger than the $Q$ of the coefficient are ignored in internal operations, which limits the scope of bit-level analysis.
+In the case of OpenFHE, we consider the following implementation details:
+  • Use of 64-bit integer coefficients (uint64_t) without arbitrary-precision arithmetic.
+  • The client can configure the initial modulus $q_{0}$.
 This parameter determines the starting bit-level precision of the scheme.
 The subsequent moduli $q_{i}$ are automatically selected to match the bit length of $q_{0}$, ensuring uniform precision across levels.
-  • The total number of moduli depends on the required multiplicative depth of the computation.
-  • Public APIs for accessing and modifying coefficients at each pipeline stage (i.e., encoding, encryption, etc.).
-  • Enables injection of fixed-length bit-flips (up to 64 bits) at controlled and reproducible positions.
-  • Supports 128-bit arithmetic, which–though not used in this work–offers a pathway for future experiments.
-IV-B Using OpenFHE as Vanilla CKKS implementation
-OpenFHE is significantly more computationally efficient than HEAAN.
-However, it does not natively offer a mechanism to disable RNS and NTT optimizations.
-To address this, we devised a procedure to emulate a vanilla CKKS scheme within OpenFHE by explicitly bypassing both RNS and NTT optimizations, despite their built-in support in the library:
+  • The plaintext and the ciphertext belong to the same ring. They both have polynomial coefficients of size $q_{0}$, regardless of the scaling factor $\Delta$.
+  • The total number of moduli depends on the required multiplicative depth of the computation. One moduli for each RNS limb.
+  • 128-bit arithmetic support, which–though not used in this work–offers a pathway for future experiments.
+IV-C2 Using OpenFHE as Vanilla CKKS Implementation
+OpenFHE is substantially more computationally efficient than HEAAN.
+Consequently, we primarily use OpenFHE in this work, although selected experiments are replicated in HEAAN for validation purposes.
+However, it does not natively offer a mechanism to disable RNS and NTT optimizations to enable OpenFHE on vanilla mode.
+We explicitly bypass both RNS and NTT optimizations.
 RNS: We enforced a single-limb representation, which is feasible for depth-zero operations (i.e., no ciphertext multiplications).
 Additionally, automatic rescaling was disabled and replaced with manual scaling to prevent the introduction of additional limbs.
-  • NTT: Prior to injecting a bit-flip, the polynomial is transformed from evaluation space back to coefficient space via inverse NTT (INTT).
+NTT: Before injecting a bit flip, we move the polynomial from evaluation space back to coefficient space using the inverse NTT (INTT).
 We flip the desired bit and transform the polynomial back to the evaluation space using NTT.
 After this hook, the pipeline continue normally.
-  • Limitations of disabling NTT: Since the NTT operates modulo $q_{0}$, flipping bits beyond the bit-width of $q_{0}$ causes modular wraparound, which can alter the intended effect of the bit-flip and potentially lead to catastrophic decryption errors.
-To minimize distortion in the analysis, we set $q_{0}$ to the maximum value supported by the library–60 bits–ensuring that most bit-flips occur within the valid range and reducing unintended side effects from modular arithmetic.
-IV-C Internal Error Detection in OpenFHE
-OpenFHE includes an internal consistency check that estimates the standard deviation of the imaginary part of the decoded plaintext, designed to detect potential secret key leakage attacks.
-Flipping high-order bits at different stages of the homomorphic pipeline can result in large deviations, triggering this mechanism and aborting the decoding.
-To isolate the analysis of numerical resilience, we disable this mechanism.
-This methodology enables controlled and reproducible exploration of both vanilla and optimized CKKS configurations, allowing for a rigorous assessment of the impact of single-bit faults at various stages of the homomorphic pipeline.
-IV-D Vanilla CKKS over OpenFHE Validation
-To ensure that OpenFHE’s vanilla mode faithfully replicates the behavior of HEAAN 1.0, we following these validation steps:
-  1. 1.
-    Configure OpenFHE in vanilla mode (not using RNS and NTT) and set $q$ and $\Delta$ to match HEAAN’s effective bit precision.
-  2. 2.
-    Restrict input values in HEAAN to real numbers, in line with OpenFHE’s plaintext constraints.
-  3. 3.
-    Compare $L_{2}$ norms and error statistics after identical bit-flips in both libraries.
-  4. 4.
-    Verify consistent sensitivity patterns across low- and high-order bit positions.
-IV-E Input Vector Generation and PRNG Control
-Since our analysis is not tied to any specific application, the input data consists of real-valued vectors with uniformly distributed random entries.
-To evaluate the sensitivity of CKKS, we define input ranges using power-of-two boundaries and sample values uniformly within these intervals.
-A wider input range affects the signal-to-noise ratio by increasing the difference between the smallest and largest representable values.
-Unless otherwise specified, input values are sampled as uniformly distributed real numbers within the interval 0 to 2560256start_ARG 0 end_ARG to start_ARG 256 end_ARG, and represented in double-precision floating-point format.
-The bounded input range allows us to test multiple $\Delta$, allowing for a more comprehensive analysis of their impact on encoding and rounding errors.
-By varying $\Delta$ across a wide range without exceeding coefficient bounds, we can better understand the trade-offs between precision and noise growth, as discussed in Section II-B.
-To evaluate stability with respect to seed variation, each experiment (i.e., each individual bit-flip) is repeated 2500 times.
-Each repetition involves 100 different seeds for the scheme PRNG and 25 for the input generator.
-These seed combinations are fixed and shared across all experiments, ensuring that results could be compared under equivalent input conditions.
-Although this number of repetitions may not suffice for exhaustive statistical inference, our tests show very low variability, confirming it is enough for the aims of this study.
-IV-F Bit-Flip Injection Procedure
-Error injection is performed by flipping a single bit in one encrypted coefficient during each execution. The general workflow for the experiments is as follows:
+Since the NTT operates modulo $q_{0}$, flipping bits beyond the bit width of $q_{0}$ causes modular wraparound, which can alter the intended effect of the bit flip and potentially lead to catastrophic decryption errors.
+To minimize distortion in the analysis, we set $q_{0}$ to the maximum value supported by the library–60 bits–ensuring that most bit flips occur within the valid range and reducing unintended side effects from modular arithmetic.
+IV-C3 Validation of Vanilla Mode using OpenFHE
+To ensure that vanilla OpenFHE mode faithfully replicates the behavior of HEAAN 1.0, we follow these validation steps, 1) Configure OpenFHE on vanilla mode (disabling RNS and NTT) and set $q_{0}$ and $\Delta$ to match HEAAN’s effective bit precision, 2) Restrict input values in HEAAN to real numbers, in line with OpenFHE’s input constraints, 3) Compare $L_{2}$-norms and error statistics after identical bit flips are introduced in both libraries and 4) Verify consistent sensitivity patterns across low- and high-order bit positions.
+IV-D Experimental Procedure
+In OpenFHE, ciphertexts and plaintexts are internally represented using NTT transformation (evaluation domain).
+Since we need to inject faults in the coefficients of the polynomials, we must ensure that we are working in the coefficient domain.
+For the vanilla and RNS modes, we first apply the inverse NTT to bring the polynomial back to the coefficient domain, then we perform the bit flip, and finally we reapply the NTT to restore the polynomial to the evaluation domain.
+For each bit flip experiment, we follow these steps:
   1. 1.
@@
   2. 2.
-    (If required) Apply the inverse NTT (INTT) to bring the plaintext or ciphertext into coefficient space.
+    Apply the inverse NTT (INTT) to bring the polynomial into coefficient space.
   3. 3.
@@
   4. 4.
-    (If required) Reapply NTT to return to evaluation space.
+    Reapply NTT to return the polynomial to evaluation space.
   5. 5.
@@
     Compute evaluation metrics.
-This process is performed once per bit and coefficient across both plaintext and ciphertext scenarios in separate experiments. The entire procedure was repeated for each seed combination.
-In OpenFHE, we test configurations both with and without RNS and NTT by dynamically disabling these optimizations at runtime.
-This allows direct comparisons between the optimized and vanilla implementations, isolating the impact of each optimization technique.
-IV-G Evaluation Metrics
-Given the original decrypted vector $\mathbf{x}$ (without bit-flips) and the resulting vector $\mathbf{y}$ after error injection and decryption, we quantify the deviation introduced by each bit-flip by two complementary evaluation metrics:
+We repeat this process for every bit position in each coefficient of the different stages of the CKKS pipeline.
+Input Vector Generation and PRNG Control
+Our analysis is application-agnostic, employing input vectors of random real numbers drawn from a uniform distribution.
+To assess the sensitivity of the CKKS scheme, we sample inputs from various intervals whose boundaries are defined by powers of two.
+Unless otherwise specified, input values are sampled as uniformly distributed real numbers within the interval 0 to 2560256, and represented in double-precision floating-point format.
+The bounded input range enables testing multiple scaling factors ($\Delta$), allowing for a more comprehensive analysis of their impact on encoding and rounding errors.
+We analyze the trade-off between precision and noise growth by varying $\Delta$ across a wide range, ensuring it stays within the coefficient bounds detailed in Section II-B.
+To evaluate stability impact of the seed variation, we repeat each experiment (i.e., each bit flip) 2500 times.
+Each repetition involves 100 different seeds for the scheme’s PRNG and 25 for the input generator.
+These seed combinations are fixed and shared across all experiments, ensuring a fair comparison using equivalent input conditions.
+The variability observed across repetitions is negligible, with no discernible differences in the resulting patterns, indicating that the adopted sampling strategy is sufficient for the purposes of this study.
+Number of Coefficients
+In order to ensure reasonable computation times given the number of seeds, we first studied the effect of the ring dimension $N$ on bit-flip sensitivity patterns.
+This parameter has the greatest impact on computational performance: a larger $N$ provides better security, but at the cost of increased computation.
+Since the bit-flip sensitivity patterns were found to be consistent across experiments with varying ring dimensions $N$ (from $2^{3}$ up to $2^{16}$), we restrict our analysis to smaller values of $N$.
+This choice both improves performance and facilitates visualization: it is far easier to discern structure in the behavior of a small set of coefficients in a single plot than to interpret patterns spread across thousands simultaneously.
+Analyzed Stages
+As presented in Section III, we corroborate experimentally that flipping a single bit in the plaintext during the Encode or Decode stages has the same effect as flipping a bit in the first polynomial ($c_{0}$) of the ciphertext during Encrypt or Decrypt.
+For this reason, we focus our study on the behavior of the ciphertext under bit-flip faults.
+This approach allows us, within a single experiment, to analyze both the behavior of the first polynomial (and thus the plaintext) and the behavior of the second polynomial of the ciphertext.
+IV-E Evaluation Metrics
+Given the original decrypted vector $\mathbf{x}$ (without bit flips) and the resulting vector $\mathbf{y}$ after error injection and decryption, we quantify the deviation introduced by each bit flip by two complementary evaluation metrics:
 Mean Squared Error (MSE)
-The element-wise error is defined as:
-$\delta_{i}=y_{i}-x_{i}$
-and the MSE is computed as:
-$\mathrm{MSE}=\sqrt{\frac{1}{k}\sum_{i=1}^{k}\delta_{i}^{2}}$
+We define the element-wise error as $\delta_{i}=y_{i}-x_{i}$, and the MSE is computed as:
+$\mathrm{MSE}=\sqrt{\frac{1}{k}\sum_{i=1}^{k}\delta_{i}^{2}}$ (18)
 Element-wise Relative Error
-For each decrypted element, we compute the relative error:
-$\epsilon_{i}=\frac{|y_{i}-x_{i}|}{|x_{i}|}$
-Results are classified into four categories based on the relative error $\epsilon_{i}$, with each category corresponding to a different tolerance level, which may vary depending on the intended application of FHE:
-  1. 1.
-    Severe Error ($\epsilon_{i}>10$): the result differs by more than one order of magnitude from the correct value, often rendering it unusable.
-  2. 2.
-    Significant Error ($0.1<\epsilon_{i}\leq 10$): the result may be unacceptable for precision tasks, but tolerable in coarse-grained applications such as visual representations or robust classification.
-  3. 3.
-    Moderate Error ($0.01<\epsilon_{i}\leq 0.1$): the deviation is small and potentially acceptable in tasks like statistical analysis or machine learning where high fidelity is not required.
-  4. 4.
-    Low Error ($\epsilon_{i}\leq 0.01$): the result is nearly indistinguishable from the correct value and is considered acceptably decrypted for most practical purposes.
-V Experimental Results
-This section presents the results obtained by injecting individual bit-flips during encryption in the CKKS scheme, as described in Section IV, with the goal of evaluating the scheme’s resilience against hardware- or software-induced faults.
-Experimental campaigns were carried out using both HEAAN 1.0 and OpenFHE 1.2.3, in their vanilla and optimized configurations (with RNS and NTT).
-V-A Vanilla CKKS
-V-A1 Evolution of the $L_{2}$ norm under Single Bit‑Flips
-To deepen the comparison, we repeated the experiments with HEAAN 1.0, obtaining results that match our theoretical predictions closely.
-This direct comparison between HEAAN and OpenFHE revealed a marked difference between the vanilla implementation and the fully optimized variant with RNS and NTT: the latter exhibits noticeably greater sensitivity to bit-flips, as evidenced by higher $L_{2}$ error magnitudes across most bit positions..
-To isolate the effect of these optimizations, we therefore configured OpenFHE in a simulated vanilla mode by disabling both NTT and RNS, as described in Sec.
-IV-B.
-In order to validate this simulation, we ran both HEAAN and OpenFHE (vanilla) under identical conditions.
-Because OpenFHE only accepts real inputs, we also restricted HEAAN accordingly to real inputs and, after decryption, considered only the real part of each result.
-This ensured a homogeneous comparison between the two libraries and confirmed that OpenFHE (vanilla) faithfully reproduces the behavior of HEAAN 1.0, in both trend and magnitude.
-Figure 7 shows the case of OpenFHE simulated as CKKS vanilla.
-Figure 7: Impact of a single-bit flip during encryption in OpenFHE without RNS or NTT.
-The horizontal axis indicates the position of the flipped bit within the ciphertext, while the vertical axis shows the resulting $L_{2}$ norm of the decoding error.
-Parameters: 60‑bit modulus, scaling factor $\Delta=2^{50}$, and ring dimension $N=4$.
-We observe that, for the first polynomial $c_{0}$, the behavior mirrors the theoretical analysis in Fig.
-4: the least significant bits exhibit higher robustness against bit‑flips, followed by an exponential increase in error magnitude.
-In the case of OpenFHE, the expected absolute robustness at coefficient $N/2$ does not hold.
-During decoding, the coefficient vector prior to the DFT (a specialized FFT) is injected with Gaussian noise affecting both its real and imaginary components.
-This noise injection is a security measure to thwart key‑recovery attacks.
-The introduced Gaussian noise has a standard deviation corresponding to the dispersion of the imaginary part of the vector being transformed, which—by properties of the scheme—also follows a Gaussian distribution.
-Flipping any bit of any coefficient, including the $N/2$ coefficient, increases this noise standard deviation, especially when higher‑order bits are modified, leading to increased noise variance that propagates across all elements of the vector, significantly altering the final output and negating the robustness at $N/2$.
-We experimentally confirmed that, when this noise injection is disabled, the observed behavior aligns with theoretical expectations: the coefficient $N/2$ of $c_{0}$ becomes robust to bit‑flips.
-Conversely, since HEAAN 1.0 does not perform this noise injection, its sensitivity matches the theory exactly.“‘
-V-A2 Impact of the Scaling Factor on Error Categorization
-One of the fundamental parameters of the CKKS scheme is the scaling factor $\Delta$.
-As analyzed in Sec.
-III, increasing $\Delta$ reduces the relative impact of a single‐bit flip during encoding or encryption and also increases the proportion of bits that remain resilient to such faults.
-To investigate this effect in practice, we repeated the previous single‐bit‐flip analysis using four different scaling factors, namely $\Delta\in\{2^{20},2^{30},2^{40},2^{50}\}$.
-Figure 8 shows the $L_{2}$ norm of the decoding error for a bit flip in the first ciphertext polynomial $c_{0}$ when running OpenFHE in simulated vanilla mode (i.e., with RNS and NTT disabled).
-For clarity, only results for $c_{0}$ are shown, since the behavior for $c_{1}$ is analogous.
-Figure 8: Impact of a single‐bit flip during encryption in OpenFHE without RNS or NTT.
-The x‑axis indicates the bit position in ciphertext $c_{0}$, while the y‑axis shows the resulting $L_{2}$ norm of the decoding error on a symmetric log scale.
-Parameters: 60‑bit modulus, scaling factors $\Delta=2^{20},2^{30},2^{40},2^{50}$, and ring dimension $N=8$.
-We observe that higher values of $\Delta$ yield lower $L_{2}$ norms after a bit flip, indicating stronger attenuation of induced errors.
-Furthermore, the point at which the error norm begins its exponential growth shifts in alignment with the magnitude of $\Delta$, confirming that larger scaling factors increase the number of error-resilient bit in each coefficient’s representation.
+For each decrypted element $i$, we compute the relative error:
+$\epsilon_{i}=\frac{|y_{i}-x_{i}|}{|x_{i}|}=\frac{|\delta_{i}|}{|x_{i}|}$ (19)
+To determine whether a vector is correctly recovered, we use the metric in Eq.
+(19) and proceed element-wise.
+For each element $i$, we compute the relative error with the reference value (i.e, same element of the vector recovered wihout no bit flip).
+We then classify the outcome of the procedure according to the fraction of elements correctly recovered:
 […]
@@
 […]
+IV-F Experimental Setup
+All experiments are execute on a dedicated workstation equipped with an Intel Core i7-11700 CPU (8 physical cores, 16 threads, 16 MB L3 cache, base 2.50 GHz) and 32 GB DDR4 RAM (dual channel, 3200 MHz).
+The operating system is Arch Linux (kernel Linux 6.9.7-arch1-1, distribution snapshot 257.5-1).
+We compiled both OpenFHE 1.3.0 and HEAAN 1.0 in release mode using g++ 15.2.1 with flags -O3.
+For HEAAN, NTL (NTL 11.5.1) was linked dynamically; OpenFHE was built with its default backend (no GPU acceleration).
+V Experimental Results
+This section presents the results obtained from injecting single bit flips during the encryption stage of the CKKS scheme, as described in Section IV, in order to evaluate its resilience against induced faults.
+V-A Vanilla CKKS
+We first analyze the vanilla CKKS scheme using OpenFHE; i.e., without RNS and NTT optimizations.
+V-A1 Evolution of the $L_{2}$-norm Under Single Bit Flips
+To evaluate the fault sensitivity of the vanilla CKKS variant, we systematically flipped one bit at a time in the ciphertexts generated during encryption, covering all bit positions in both $c_{0}$ and $c_{1}$ polynomials.
+After each injection, the modified ciphertext was decrypted and decoded, and the $L_{2}$-norm of the difference with respect to the fault-free output was computed.
+This metric quantifies the magnitude of the decoding error associated with each bit position.
+Figure 8 shows the resulting $L_{2}$-norm distribution for all possible single-bit flips, highlighting which regions of the ciphertext are most sensitive to faults.
+The ring dimension is $N=4$, with modulus $q_{0}=60$-bit and scaling factor of $\Delta=2^{50}$.
+Figure 8: Bit flip during encryption on vanilla mode.
+The horizontal axis indicates the position of the flipped bit within the ciphertext, while the vertical axis shows the resulting $L_{2}$-norm of the decoding error.
+Parameters: 60‑bit $q_{0}$, input range from 0 to 2560256, $\Delta=2^{50}$, and $N=4$.
+We observe that, for the first polynomial $c_{0}$, the behavior closely follows the theoretical trend shown in Figure 4: least significant bits exhibit higher resilience to bit flips, while the error magnitude grows exponentially from the bit position corresponding to $\Delta$ up to the bit position defined by the modulus $q_{0}$, after which it abruptly drops to zero.
+In our theoretical analysis, as well as in HEAAN, the coefficient $N/2$ of $c_{0}$ is completely robust.
+However, in OpenFHE, as we can observe in Figure 8 near bit 192, this behavior does not occur.
+This difference arises because OpenFHE, during decoding, injects Gaussian noise into the plaintext prior to the DFT (a specialized FFT), perturbing both real and imaginary components as a countermeasure against key-recovery attacks.
+The injected noise has a standard deviation proportional to the spread of the imaginary part of the pre-DFT vector (plaintext), which, due to the scheme’s properties, follows a Gaussian distribution.
+Flipping any bit in any coefficient, including $N/2$, alters this spread; higher-order bit flips have a stronger effect, increasing the overall noise variance and propagating it across all polynomial coefficients, thereby removing the theoretical robustness of coefficient $N/2$.
+Disabling this noise injection in OpenFHE restores the expected behavior: coefficient $N/2$ of $c_{0}$ becomes robust to bit-flips.
+Since HEAAN 1.0 does not include this security noise, its experimental sensitivity matches the theoretical prediction exactly.
+V-A2 Impact of the Scaling Factor
+One of the key parameters of the CKKS scheme is the scaling factor $\Delta$.
+As analyzed in Section III, increasing $\Delta$ reduces the relative impact of a single‐bit flip during encoding or encryption and also increases the proportion of bits that remain resilient.
+To validate this behavior experimentally, we analyze the impact of three different scaling factors, namely $\Delta\in\{2^{20},2^{40},2^{50}\}$.
+Figure 9 shows the $L_{2}$-norm of the decoding error for a bit flip in the first ciphertext polynomial $c_{0}$ on vanilla mode.
+For clarity, only results for $c_{0}$ are shown, since the behavior in $c_{1}$ is analogous (as is shown in Fig.
+8).
+Figure 9: Bit flip during encryption on vanilla mode.
+The horizontal axis indicates the bit position in ciphertext $c_{0}$, while the vertical axis shows the resulting $L_{2}$-norm of the decoding error on a symmetric log scale.
+Parameters: 60‑bit $q_{0}$, $\Delta\in\{2^{20},2^{40},2^{50}\}$, input range from 0 to 2560256 and $N=8$.
+We observe that higher values of $\Delta$ yield lower $L_{2}$-norms after a bit flip, indicating stronger attenuation of induced errors.
+Furthermore, the point at which the error norm begins its exponential growth shifts in alignment with the magnitude of $\Delta$, confirming that larger scaling factors increase the number of error resilient bits in each coefficient.
+Figure 10: Histogram of error‐categorization outcomes for different scaling factors.
+Each group of bars corresponds to a specific $\Delta$, showing the percentage of trials in the different categories.
+Parameters: 60‑bit $q_{0}$, $\Delta\in\{2^{20},2^{40},2^{50}\}$, input range from 0 to 2560256 and $N=2^{13}$.
+In Fig.
+10, we adopt a more realistic configuration using a ring dimension $N=2^{13}=8192$ (4096 encryptable elements).
+For each scaling factor $\Delta\in\{2^{20},2^{40},2^{50}\}$ and each single-bit flip, we track the number of elements that decrypt correctly.
+We use the three categories defined in Sec.
+IV-E based on element-wise relative error (Eq.
+(19)).
+As $\Delta$ increases, the frequency of Robust outcomes (green bars) also increases, confirming that a larger scaling factor provides greater tolerance to faults.
+Although we do not explicitly compute the $L_{2}$-norm threshold at which decryption completely fails, the observed consistency of the Application-dependent error category across scaling factors indicates that the width of the transition interval—from full resilience to total failure—remains approximately constant for all values of $\Delta$.
+Within this transition region, some elements decrypt correctly while others fail, reflecting the gradual loss of robustness rather than an abrupt threshold.
+We observe a slight increase in the Application-dependent category for $\Delta=2^{20}$ and more markedly for smaller $\Delta$ (data not shown).
+This is consistent with the chosen parameters: with $q_{0}=$60\text{\,}\mathrm{bit}$\$ and $\Delta=2^{20}$ the scheme provides roughly 20 fractional bits and about 40 integer bits.
+When input values are limited to the range up to $2^{8}=256$, most of the available integer range remains unused, so small perturbations introduced during encoding/decoding (rounding, or bit-flip noise) can move some elements to the Application-dependent category.
+V-A3 Impact of the Input Value’s Magnitude
+We fix $\Delta=2^{20}$ and $q_{0}=$60\text{\,}\mathrm{bit}$\$ to enable a wider representable input range.
+As previously explained, the number of bits available to represent the integer part of an input value is approximately $q_{0}-\log_{2}(\Delta)$–about 40 bits in this configuration.
+In our experiments, we define input value intervals, from which we uniformly sample the elements.
+Figure 11 shows histograms of the error category distributions for three input intervals: $2^{9},2^{10}]$, $2^{19},2^{20}]$, and $2^{29},2^{30}]$.
+We deliberately avoid the interval $2^{39},2^{40}]$ because it may overflow and cause a failed decryption even with unmodified inputs (i.e., without bit flips).
+The selected intervals are representative of the range of input magnitudes considered in our experiments.
+We use the following parameters: ring dimension $N=8192$, ciphertext modulus $q_{0}=$60\text{\,}\mathrm{bit}$\$, and scaling factor $\Delta=2^{20}$.
+Figure 11: Histogram of error‐categorization outcomes for different scaling inputs.
+Each group of bars corresponds to a specific input range, showing the percentage of trials in the different categories.
+Parameters: 60‑bit $q_{0}$, $\Delta=2^{20}$, input range $\in[2^{9},2^{10}],[2^{19},2^{20}]$ and $2^{29},2^{30}]$ and $N=2^{13}$.
 […]
-For this experiment, we adopted a realistic setting with ring dimension $N=2^{13}=8192$ (i.e., 4096 encrypted slots).
-For each $\Delta\in\{2^{20},2^{30},2^{40},2^{50}\}$ and each single‐bit flip, we counted how many slots were correctly decrypted and assigned the result to one of the three categories.
-Figure 9 presents a histogram of the category frequencies for each scaling factor.
+V-A4 Effect of Reducing the Number of Encrypted Elements
+As noted previously, CKKS can encode up to $N/2$ elements, where $N$ is the power-of-two ring degree.
+When we encode fewer elements–still a power of two–we observe that additional coefficients remain unaffected by single‐bit flips.
+Figure 12 shows results for $N=16$ with only $N/4=4$ encrypted elements.
+Figure 12: Bit flip during encryption on vanilla mode.
+The horizontal axis indicates the bit position in the ciphertext, while the vertical axis shows the resulting $L_{2}$-norm of the decoding error.
+Parameters: 60‑bit $q_{0}$, $\Delta=2^{20}$, sample input range from 0 to 2560256, $N=16$, and four encrypted elements.
+The results reveal that odd‐indexed coefficients of polynomial $c_{0}$ remain unchanged under bit flips, whereas even‐indexed coefficients exhibit the same sensitivity patterns observed earlier.
+As explained in Section II, the ratio between the maximum number of encryptable elements ($N/2$) and the number actually used defines an interval (or gap) between the coefficients participating in the inverse transform.
+During the special FFT, only coefficients which indices are multiple of this gap are considered and scaled in the input vector (see Section I for details).
+For example, when encoding $N/2$ elements, all $N$ coefficients of $c_{0}$ are sensitive to bit flips, but encoding $N/4$ elements reduces the number of sensitive coefficients of $c_{0}$ to $N/2$.
+In the case of $c_{1}$, all coefficients are sensitive to bit changes.
+V-B Resilience in RNS-only mode
+We focus on the RNS-only configuration when the number of encrypted elements is reduced to half or less of the available encryptable elements.
+Figure 13 reports results for ring dimension $N=16$ with only four encrypted elements, scaling factor $\Delta=2^{20}$, and 8‐bit input values.
+Figure 13: Bit flip during encryption on RNS-only mode with two limbs.
+The horizontal axis indicates the flipped bit position; the vertical axis indicates whether decoding succeeded (robust category) or failed (catastrophic category).
+Parameters: 60‑bit $q_{0}$, input range from 0 to 2560256, $\Delta=2^{20}$, $N=16$, and four encrypted elements.
+We observe that odd‐indexed coefficients of $c_{0}$ remain fully robust: all the input elements decrypt correctly.
+However, any bit flip in an even‐indexed coefficient, causes catastrophic failures.
+The number of sensitive coefficients exactly matches the gap analysis from Section V-A4.
+Polynomial $c_{1}$ exhibits no robust bits at all: every bit flip causes decrypt failure.
+These results show that running the RNS-only mode leaving some unused elements can unveil robustness.
+To ensure a secure-enough schemes, CKKS requires a ring dimension in the order of $2^{15}$, which leads to a payload of near $2^{14}$ encrytable elements.
+In many practical uses, the amount of needed elements is significantly less, allowing a fault tolerance scheme by taking advantage of the unsed input elements.
+MNIST Example.
+To illustrate the practical impact of bit-flip resilience in CKKS, we consider a case using input images from the MNIST dataset [23].
+Each image consists of $28\times 28=784$ grayscale pixels, where each pixel is encoded using integer values ranging from 0 to 255.
+To encode 784 values using CKKS batching, we need a vector size that is a power of two and greater than or equal to 784; the smallest size is $2^{10}=1024$.
+Since CKKS requires a ring dimension $N$ that is twice the batch size, we need $N\geq 2^{11}=2048$.
+However, to comply with security requirements, we select a ring dimension of $N=2^{15}$.
+Only $2^{11}$ coefficients of the plaintext or ciphertext polynomials store actual image data, while the remaining $2^{15}-2^{11}$ coefficients are unused.
+Figure 14 shows results for $N=2^{15}$, and $2^{10}$ encrypted elements of MNIST images, $q_{0}$ modulus of 60 bits, scaling factor of $2^{20}$, and three-limb RNS.
+As expected, flipping any bit in the $c_{1}$ polynomial consistently produces an undecryptable result, whereas in $c_{0}$ only about one in every sixteen coefficients is sensitive to bit-flips.
 […]
-[…]
-V-A3 Effect of the Effective Representation Range on Error Categorization
-We now analyze the impact of the effective dynamic range of the input data representation.
-We fix $\Delta=2^{20}$, which allows for a wider input range.
-Recall that the number of bits available for the integer part is approximately given by $q_{0}-\Delta$.
-The effective dynamic range refers to the ratio between the largest and smallest values that can be represented in fixed-point encoding.
-In our experiments, this range is controlled indirectly by limiting the number of bits allocated to the fractional part of the input values.
-[…]
-[…]
-[…]
-We observe behavior similar to the impact of the scaling factor, as seen in Fig.
-9.
-As we increase the size of the input space, the percentage of correctly decrypted elements increases while the percentage of incorrectly decrypted elements decreases.
-If the size of input elements grows while maintaining a fixed scaling factor, this allows representing values with greater relative precision compared to the existing noise.
-V-A4 Effect of reducing the number of encrypted elements
-As noted previously, CKKS can encode up to $N/2$ slots, where $N$ is the ring degree (a power of two).
-When we encode fewer slots—still a power of two—we observe that additional coefficients remain unaffected by single‐bit flips.
-Figure 11 shows results for $N=16$ with only $N/4=4$ encrypted slots, allowing us to inspect a larger set of coefficients and evaluate their sensitivity.
-Figure 11: Impact of a single‐bit flip during encryption in OpenFHE without RNS or NTT.
-The horizontal axis indicates the bit position in the ciphertext, while the vertical axis shows the resulting $L_{2}$ norm of the decoding error.
-Parameters: 60‑bit modulus, scaling factor $\Delta=2^{20}$, ring dimension $N=16$, and 4 encrypted slots.
-The results reveal that odd‐indexed coefficients of polynomial $c_{0}$ remain unchanged under bit‐flips, whereas even‐indexed coefficients exhibit the same sensitivity patterns observed earlier.
-This phenomenon stems from CKKS’s decoding process in both OpenFHE and HEAAN.
-As explained in Sec.
-II, the ratio between the maximum number of slots ($N/2$) and the number actually used defines an interval—termed the gap—between the coefficients participating in the inverse transform.
-During the “special” FFT, only coefficients whose indices are multiples of this gap are considered and scaled in the input vector (see Sec.
-I for details).
-Consequently, when encoding $N/2$ slots, all $N$ coefficients of $c_{0}$ are sensitive to bit‐flips; encoding $N/4$ slots reduces the number of sensitive coefficients of $c_{0}$ to $N/2$; and so on.
-In the case of $c_{1}$, no matter the number of slots used, all coefficients are sensitive to bit changes.
-Experimental campaigns confirm that our OpenFHE vanilla simulation (with RNS and NTT disabled) matches the expected behavior.
-V-B Resilience under an intermediate RNS mode without NTT
-We next conducted experimental campaigns to assess CKKS’s sensitivity when using only one of the numeric optimizations—either NTT or RNS.
-I n both pure-NTT and pure-RNS configurations, any single‐bit flip caused decryption errors of unacceptable magnitude.
-As shown in our theoretical analysis, bit flips under RNS produce errors several orders of magnitude larger than under NTT, while NTT alone still yields extremely large errors that preclude practical precision.
-An interesting case emerges when CKKS is run with RNS only and the number of encrypted slots is halved or more.
-Figure 12 reports results for ring dimension $N=16$ with only 444 encrypted slots (i.e., one power‐of‐two below the maximum), scaling factor $\Delta=2^{20}$, and 8‐bit input values.
-[…]
-[…]
-Polynomial $c_{1}$ exhibits no robust bits at all: every bit flip causes uniform decoding failure, so its error‐categorization histogram is flat and standard deviation is zero.
-These results demonstrate that running CKKS with RNS only—together with a non‐maximal but non‐zero gap—can concentrate robustness on a subset of coefficients.
-Such a configuration suggests a trade‐off between ring utilization and fault resilience.
-Future work may explore tailored RNS and NTT parameterizations to minimize sensitivity to bit flips while preserving sufficient slot capacity, laying the groundwork for error‐reduction strategies in homomorphic encryption schemes that balance precision and fault tolerance.“‘
-V-C Summary and Comparison
-We provides a high‑level comparison of the different CKKS configurations evaluated in this work. For each scenario, we highlight the key resilience pattern and any notable observations.
-  • Vanilla CKKS:
-    • Least‐significant bits exhibit robustness; error grows exponentially for higher bits.
+Figure 14: Bit flip during encryption on RNS-only mode with three limbs, using MNIST images as input.
+The horizontal axis indicates the flipped bit position; the vertical axis indicates the resulting L2 norm of the decoding error.
+Only $c_{0}$ is shown because $c_{1}$ is all undecryptable similarly as Fig.
+13.
+Parameters: 60-bit $q_{0}$, $\Delta=2^{20}$, $N=2^{15}$, and $2^{10}$ encrypted elements.
+However, this resiliency comes at a substantial computational cost.
+As explained in Section II, NTT-based multiplication enables fast polynomial product computations with a complexity of $\mathcal{O}(N\log N)$, where $N$ is the ring dimension.
+When NTT is disabled, the underlying polynomial multiplication reverts to a schoolbook approach with complexity $\mathcal{O}(N^{2})$.
+As a result, homomorphic operations become orders of magnitude slower when operating in coefficient space without NTT.
+Although we do not perform direct runtime benchmarking in this study, prior works such as [6] report performance slowdowns of approximately ten times when switching from the RNS+NTT-optimized scheme to the vanilla CKKS implementation.
+In scenarios where RNS is used but NTT is disabled, the degradation is expected to be smaller, since RNS still reduces the cost of large-integer arithmetic by enabling parallel residue computations.
+In practice, the slowdown depends on various factors, including the specific implementation, the chosen ring dimension, and the availability of hardware-level optimizations (e.g., SIMD instructions or parallel FFTs).
+This highlights a fundamental trade-off between fault resilience and computational efficiency, which must be carefully balanced in real-world deployments.
+VI Conclusions
+This paper presents a comprehensive evaluation of CKKS’s sensitivity to single‐bit flips, emphasizing the influence of RNS and NTT optimizations on error resilience.
+While the vanilla CKKS implementation exhibits substantial inherent tolerance to individual bit‐flip faults, the optimized variants–with RNS or NTT enabled–show markedly increased vulnerability, potentially undermining the correctness of decrypted outputs.
+We provide a high‑level comparison of the different CKKS configurations evaluated in this work. For each scenario, we highlight the key resilience pattern and any notable observations.
+Vanilla CKKS
+Least‐significant bits exhibit robustness; error grows exponentially for higher bits up to modulus $q_{0}$.
 Coefficient $N/2$ robust in HEAAN 1.0 but not in OpenFHE vanilla (due to Gaussian noise injection).
-  • Varying $\Delta$:
-    • Larger $\Delta$ delays the onset of exponential error growth to more significant bits.
-    • $\Delta\geq 2^{30}$ yields $\mathbf{Tail}$ outcomes in over 80% of trials.
-  • Effective Representation Range:
-    • Increasing integer‐part range raises $\mathbf{Tail}$ frequency and reduces $\mathbf{Head}$ frequency.
-    • Behavior parallels that of increasing $\Delta$.
-  • Reduced Slots:
-    • Encoding fewer slots reduce the number of sensitive coefficients in $c_{0}$.
-  • RNS without NTT:
-    • Some coefficients of $c_{0}$ remain robust; all flips in $c_{1}$ fail.
-    • Demonstrates a “gap”‐driven concentration of robustness when using RNS only.
-This concise overview highlights how each parameter adjustment—from vanilla through optimized variants—modulates CKKS’s fault tolerance, guiding practical parameter selection under two‐column constraints.
-VI Conclusions
-This paper has presented a comprehensive evaluation of CKKS’s sensitivity to single‐bit flips, emphasizing the influence of RNS and NTT optimizations on error resilience.
-While the vanilla CKKS implementation exhibits substantial innate tolerance to individual bit‐flip faults, the optimized variants—with RNS or NTT enabled—display markedly increased vulnerability, potentially undermining the correctness of decrypted outputs.
+Varying $\Delta$
+Larger $\Delta$ shifts the onset of exponential error growth towards more significant bits. $\Delta\geq 2^{30}$ yields robust outcomes in over 80% of trials.
+Effective Representation Range
+Increasing integer‐part range raises catastrophic frequency and reduces robust frequency. Behavior mimics that of increasing $\Delta$.
+Fault tolerance based on unused input elements
+Encoding fewer elements increases the number of robust coefficients in $c_{0}$ or plaintext. Not using NTT increases substantially the computational cost.
 Our findings indicate that, despite its performance drawbacks, vanilla CKKS offers a robustness advantage by minimizing the impact of isolated bit errors compared to its optimized counterparts.
-Furthermore, when employing large ring dimensions for security, encoding fewer slots (powers of two below the maximum) introduces redundancy that renders a subset of coefficients entirely insensitive to bit flips.
-This “gap”‐driven effect persists even in optimized configurations, highlighting a practical mechanism for enhancing resilience without altering core codec operations.
-Overall, these results deepen our understanding of fault tolerance in homomorphic encryption schemes and provide actionable guidance for balancing throughput, security, and error resilience when selecting CKKS parameters.
-Future Work
-Promising directions include studying error tolerance in application‐specific domains—such as deep neural networks for image processing, where limited accuracy loss is acceptable—and conducting a detailed investigation of error propagation under multiple concurrent bit flips.
-Such analyses will further illuminate strategies for error mitigation in practical homomorphic computing scenarios.
+Furthermore, when employing large ring dimensions for security, encoding fewer elements (powers of two below the maximum) introduces redundancy that renders a subset of coefficients robust to bit-flips.
+This gap‐driven effect persists even in some optimized configurations, highlighting a practical mechanism for enhancing resilience at the cost of performance.
+These results provide actionable guidance for balancing throughput, security, and error resilience when selecting CKKS parameters.
+They also lay the groundwork for future research in the context of application-specific error tolerance–e.g., deep neural networks for image processing–where limited accuracy loss is acceptable, as well as evaluation in multi-bit error scenarios.
 Acknowledgments
-First, I would like to thank Pradip Bose, whose inspiration to explore fault tolerance in homomorphic encryption significantly shaped this work.
-I also appreciate the open‐source homomorphic encryption community and the developers of the OpenFHE and SEAL libraries, whose tools and documentation greatly facilitated our experiments.
-In particular, I extend my gratitude to Ahmad Al Badawi for his help setting up OpenFHE correctly and for his clear explanations and references on CKKS encoding and decoding.
+We acknowledge the support from the open‐source homomorphic encryption community and the developers of the OpenFHE and HEAAN libraries.
+We extend our gratitude to Dr.
+Ahmad Al Badawi for his help setting up OpenFHE and clear explanations.
+This work is partially supported by grants from Universidad de Buenos Aires (UBA) UBACYT 20620190100001BA, and CONICET PIP 11220220100027CO.